### v1

# suppReferences

Abstract
Misinformation undermines public trust in science and democracy, particularly on social media where inaccuracies can spread rapidly. Experts and laypeople have shown to be effective in correcting misinformation by manually identifying and explaining inaccuracies. Nevertheless, this approach is difficult to scale, a concern as technologies like large language models (LLMs) make misinformation easier to produce. LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce plausible but false content and references, and limitations in addressing multimodal information. To address these issues, we propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information. By retrieving contextual evidence and refutations, Muse can provide accurate and trustworthy explanations and references. It also describes visuals and conducts multimodal searches for correcting multimodal misinformation. We recruit fact-checking and journalism experts to evaluate corrections to real social media posts across 13 dimensions, ranging from the factuality of explanation to the relevance of references. The results demonstrate Muse’s ability to correct misinformation promptly after appearing on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality corrections from laypeople by 29%. This work underscores the potential of LLMs to combat real-world misinformation effectively and efficiently.
Introduction
Misinformation—including misinformed, inaccurate, and misleading information—has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5]. It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7]. Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S. Capitol, where five people died and more than 100 police officers were injured[8]. Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10]. The result is tragic; an analysis showed that COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S. alone[11]. Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
Fortunately, high-quality and prompt misinformation correction, which broadly includes identifying what part(s) of content is (or is not) misinformation and explaining why that part of the content is misinformed, inaccurate, or misleading, has effectively reduced people’s false beliefs[17, 4, 18]. To this end, domain experts (e.g., on FactCheck.org) have been instrumental in providing accurate, detailed, and well-written explanations of whether and why a claim can be inaccurate with references. Subsequently, other research has shown that groups of laypeople can also identify and produce effective explanations to correct misinformation if their assessments are strategically aggregated[18, 19, 20]. Nevertheless, combating misinformation in the digital and generative artificial intelligence (AI) era cannot solely rely on human intelligence[21, 22]. The number of active fact-checking sites has remained stagnant since 2022[23]. Even though crowdsourcing is more scalable than expert-led activities, keeping pace with millions of posts daily on social media is impossible. As Brandolini’s law indicates[24], correcting misinformation is laborious—often requiring one to search and evaluate the credibility of related articles with evidence completing or refuting the claims in the given piece and write responses as justifications—whereas creating and spreading misinformation is easy. The absence of sufficient coverage markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[25].
While generative AI models such as LLMs (e.g., OpenAI’s GPT series[26]) raise concerns that they facilitate the creation of misinformation, they also potentially make scaling up the correction of misinformation possible. Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand. Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[27, 28, 29]. However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[30]. For instance, GPT-4[26] (as of March 2023) and MisinfoCorrect[31] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics. They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[30, 32]. A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39, 40]. However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating inaccuracies and backfiring (i.e., reinforcing rather than reducing false beliefs)[34, 41, 42]. These LLMs, as well as MisinfoCorrect[31], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs. Finally, understanding the quality of a correction necessitates a comprehensive human evaluation[30]. However, prior uses of LLMs for misinformation correction were evaluated by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[31])[40], which is insufficient to comprehensively understand their performance in identifying and explaining inaccuracies, generated text, and references.
In this article, we propose Muse, a scalable approach for multimodal misinformation correction. Muse makes use of an LLM and augments it with the ability to handle visuals, access timely, factual, and credible knowledge on the web, retrieve evidence that refutes or contextualizes misinformation, and generate clear explanations with accurate and trustworthy references. We recruited experts in fact-checking and journalism to comprehensively evaluate corrections made by Muse, and compare them to baselines including GPT-4 and high-quality corrections written by the collective efforts of laypeople on social media. Our assessment measures the overall quality of corrections, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness of corrections, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references. We find that overall Muse outperforms GPT-4 by 37% and high-quality corrections from laypeople by 29%, demonstrating Muse’s ability to correct both textual and visual misinformation across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine, promptly after appearing on social media.
Approach
Muse is designed to automatically generate a correction to a given piece of social media content that is potentially misinformation. As the input, the content can contain text with or without visual information. As the output, the correction should identify what part(s) of the content is (or is not) misinformation, explain why that part of the content is misinformation, and provide links as references. The general pipeline of Muse is illustrated in Fig. 1.
We start by introducing the details of Muse with a piece of text-only misinformation as the input. First, Muse generates queries based on an LLM from the misinformation (Fig. 1b; Methods). Each query acts as the input of a web search engine to access timely updated web content and obtain a list of web links directly relevant to the query (Fig. 1b; Methods). After scraping the content from these web links, Muse calculates their direct relevance to the misinformation and removes irrelevant web pages (Fig. 1b; Methods). Then, Muse determines the credibility of web pages by looking up their publishers’ factuality and bias ratings and selects pages with high factuality and minimal bias (Fig. 1c; Methods). Next, Muse leverages an LLM to extract sentences or paragraphs from each of the web pages as evidence, i.e., the text either explicitly refutes the misinformation or provides additional context (Fig. 1d; Methods). Finally, Muse generates a correction to the misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig. 1d; Methods).
Note that misinformation, especially when it comes to social media, often contains extraneous information that is irrelevant to a correction, including unverifiable opinions or emojis, such as the textual content of the misinformation in Fig. 1. Therefore, generating queries instead of simply using the text of the post improves web searches as a way of denoising misinformation; see example queries in Fig. 1b. Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs to be corrected, whereas generating one query may overlook some misinformed claims (Supplementary Fig. S1). Another concern may arise from filtering retrieved web pages based on how relevant the content is to the misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction. However, the increase in selected web pages drastically elevates the expense of Muse (Methods). We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating corrections (Supplementary Fig. S2). Moreover, as illustrated in Fig. 1c, Muse ranks the selected web pages by their publishers’ factuality and bias. It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation).
When the input is a piece of multimodal misinformation with textual and visual information, Muse first describes each image in text (Fig. 1a) so that the misinformation can be handled by any LLM on downstream tasks, including query generation (Fig. 1b), evidence extraction (Fig. 1d), and correction generation (Fig. 1d). Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig. 1a; Methods). Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate corrections[43]. For example, even a state-of-the-art image captioning model[44] may describe the visual misinformation in Fig. 1 as simply “list of banned books in Florida”. The description overlooks the listed titles of books that are essential for the visual misinformation’s correction (see more examples in Supplementary Fig. S3). In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig. 1b; Methods).
Evaluation
Muse’s evaluation was based on X Community Notes data. Community Notes (communitynotes.twitter.com, formerly Twitter Birdwatch) was launched in 2021 to empower people on X to collaboratively add context to potentially misleading posts. It has been shown to effectively reduce the spread of misinformation[20]. Community Notes operates by enabling people on X, often laypeople, to identify whether a tweet is potentially misleading, write free-response corrections of any tweet, and assess whether other people’s corrections are helpful or unhelpful. The platform strategically aggregates people’s assessments with a transparent algorithm such that a correction has a higher helpfulness score if more people with diverse backgrounds (e.g., with different political ideologies) agree that the correction was helpful[20]. Besides canceling out inaccurate judgments made by some individuals, the algorithm has been shown to empirically address the “tyranny of the majority” and coordinated attempts at gaming[20, 19]. A correction with a sufficiently high helpfulness score is then displayed on the corresponding tweet on X and visible to the public[20] (Methods; Supplementary Fig. S4).
We selected all potentially misleading tweets from Community Notes that have at least two corrections, where one correction has a high helpfulness score and the other has an average helpfulness score (Methods). We further generated corrections to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods). Experts in fact-checking and journalism were recruited, onboarded, and asked to evaluate the quality of corrections by various approaches to the same tweet (Methods). Experts were blinded to which approach had generated each correction. The evaluation contains 13 specific criteria, covering how well the correction identifies and explains inaccuracies, the quality of generated text, and the quality of references (Methods). It also contains the overall quality of correction by taking all 13 evaluation criteria into account (Methods).
Our primary finding is that the overall quality of Muse-generated corrections is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 2a). The overall quality of Muse-generated corrections has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness corrections (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated corrections (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness corrections (mean: 5.2; $p=5\times 10^{-81}$; $N=462$). Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness corrections and GPT-4-generated corrections ($p=0.4$; $N=464$), the overall quality of GPT-4-generated corrections has the highest variability, and GPT-4 generates more corrections with extremely low quality. The standard deviation of the overall quality of GPT-4-generated corrections is 2.7, vs only 2.0 for Muse and laypeople. 10% of GPT-4’s generated corrections have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness corrections, 3% for laypeople’s high-helpfulness corrections, and 2% for Muse-generated corrections. Note that laypeople’s corrections were created on average 14 hours after the tweet was posted on social media. Here, Muse only retrieved web pages published before the tweet was posted (Methods).
Examining specific components of correction quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in identifying and explaining inaccuracies (Fig. 2b-f). Experts assessed that Muse-generated corrections more explicitly identify and explain where and why the tweets are inaccurate than GPT-4’s and laypeople’s high-helpfulness corrections (Fig. 2b). 89% of Muse’s generated corrections explicitly identify and explain inaccuracies, 16% more than GPT-4-generated corrections, 29% more than laypeople’s high-helpfulness corrections, and 43% more than laypeople’s average-helpfulness corrections (Fig. 2b). As for identifying where the tweets are inaccurate, we found that Muse more comprehensively identifies the tweets’ inaccuracies with fewer mistakes (i.e., claiming where the tweets should be inaccurate as accurate or vice versa) than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2c-d). 91% of Muse’s generated corrections have at least one accurate identification without any mistake, 11% more than laypeople’s high-helpfulness corrections, 19% more than GPT-4-generated corrections, and 26% more than laypeople’s average-helpfulness corrections (Fig. 2c). Muse has 61% of generated corrections accurately identifying all the tweets’ inaccuracies, vs GPT-4 has 38%, laypeople who produce high-helpfulness corrections have 26%, and laypeople who produce average-helpfulness correction have 17% only (Fig. 2d). Furthermore, Muse explains why the tweets are inaccurate more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig. 2e-f). 70% of corrections by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness corrections, 47% by GPT-4, and 37% for laypeople’s average-helpfulness corrections only (Fig. 2e). Meanwhile, the average informativeness score of Muse-generated corrections is 8.4, 18% higher than GPT-4-generated corrections, 29% higher than laypeople’s high-helpfulness corrections, and 50% higher than laypeople’s average-helpfulness corrections (Fig. 2f).
Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of generated text (Fig. 2g-k). Muse, which can augment GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals, exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text in correcting misinformation compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$). Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness corrections by laypeople and additionally less toxic than the text of average-helpfulness corrections by laypeople ($p=4\times 10^{-12}$; $N=462$). In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness corrections by laypeople, and 43% higher than the text of average-helpfulness corrections by laypeople (Fig. 2g). 74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness corrections by laypeople and 45% for GPT-4-generated text (Fig. 2h). Muse has 99% generated text without any mistake in the use of English, vs laypeople who produce even high-helpfulness corrections have 89% such text (Fig. 2i). Around 91% of Muse-generated text is highly coherent and logical, vs 76% and 61% for the text of high- and average-helpfulness corrections by laypeople (Fig. 2j). Around 99% of Muse-generated text is not biased, impolite, and provoking, 12% higher than that of average-helpfulness corrections by laypeople (Fig. 2k).
Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of references (Fig. 2l-n). First, GPT-4 hallucinates references frequently. 49% of its links result in “page-not-found” errors (Fig. 2l), and only 76% of reachable links support or are relevant to the generated text (Fig. 2m). Muse significantly reduces such hallucinations with nearly 100% links being reachable (Fig. 2l) and 96% reachable links being relevant to the generated text (Fig. 2m). Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness corrections by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig. 2n).
Furthermore, we observed that the quality of Muse-generated corrections to text-only and multimodal (textual and visual) misinformation is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig. 3a; Supplementary Fig. S5-S6). The quality of Muse-generated corrections to text-only tweets is 29% higher than GPT-4-generated corrections ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness corrections ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness corrections ($p=10^{-67}$; $N=310$). The quality of Muse-generated corrections to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness corrections ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness corrections ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
Finally, we found that the quality of Muse-generated corrections to misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Methods; Fig. 3b; Supplementary Fig. S7-S11). The quality of Muse-generated corrections is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated corrections. The quality of Muse-generated corrections is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness corrections by laypeople.
Discussion
While concerns have arisen about LLMs in facilitating the creation of misinformation[21, 22], our work demonstrates the potential of LLMs to improve the online information ecosystem by correcting misinformed, inaccurate, and misleading information[2, 3]. Correcting misinformation, particularly on social media, is complex. It requires comprehending multimodal content, understanding context involving emerging topics or events, accurately and comprehensively identifying stated inaccuracies that are subtly different from the facts, and writing easy-to-understand, precise, and informative explanations supported by trustworthy references. We propose Muse and demonstrate the high quality of its automatically generated corrections to misinformation on social media. Muse augments existing powerful LLMs (e.g., GPT-4[26]) with the capabilities of addressing visuals, accessing up-to-date knowledge, and finding accurate references nonparametrically (i.e., without the need to train or fine-tune a machine learning model) such that Muse can be easily and cheaply updated, especially compared to parametric models[35]. Results further validate that addressing visual information is crucial in correcting multimodal misinformation, where GPT-4 struggles but Muse excels (Fig. 3a). Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining inaccuracies (Fig. 2c,e), generating text that is more factual and relevant to the responded content (Fig. 2g-h), and providing more references that are real and relevant to the generated text (Fig. 2l-m) than GPT-4.
We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently correcting misinformation. Muse is end-to-end and thereby simple to use. Meanwhile, our results reveal that Muse-generated corrections have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness corrections in correcting misinformation across modalities and domains (Fig. 2-3). Besides the highest accuracy and factuality (Fig. 2c,e,g-h,l-m), Muse’s generated corrections show the highest readability by being the most explicit, fluent, and coherent (Fig. 2b,i-j). Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating corrections with the least toxic text and most credible references[31, 41, 42] (Fig. 2k,n). It can also reduce the implied truth effect (i.e., increasing people’s perceived accuracy of overlooked inaccuracies)[25] by comprehensively identifying all the inaccuracies in a social media post (Fig. 2d) and being more capable of correcting misinformation at scale. These high-quality corrections by Muse can generally be obtained two minutes after misinformation appears on social media (Methods). By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify corrections themselves (Fig 2,3f).
Muse’s corrections cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods). Our focus in designing Muse was in maximizing the quality of corrections. Considering the complexity of this task and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives. For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without explaining the inaccuracies.
This work also faces the following limitations. First, although Muse is capable of addressing and correcting multimodal misinformation with text and images, it cannot accept video inputs. Second, we focus on correcting misinformation in a single language, English, one of the most spoken languages in the world, whereas misinformation has become a global issue and exists in various high- and low-resource languages [45]. Third, we evaluated Muse using real-world misinformation on a single social media platform, X, as its Community Notes system provides publicly accessible large-scale data and a transparent state-of-the-art aggregation algorithm, enabling us to compare Muse’s corrections with the high- and average-helpfulness corrections made by social media users. X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13]. Fourth, experts assessed and compared Muse against one other LLM, GPT-4[26], which can also be seen as an ablation study, since Muse augments GPT-4 (Methods). GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
Conclusion
We proposed Muse, a nonparametric LLM for scaling up high-quality and prompt corrections of misinformation on social media, that can integrate textual and visual information, access timely updated knowledge, and generate easy-to-understand corrections with accurate and trustworthy references. Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness corrections on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of corrections overall. Muse also excels in correcting text and multimodal (textual and visual) misinformation and misinformation across politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine. This study demonstrates the potential of LLMs in correcting misinformation online effectively, scalably, promptly, and transparently.
Acknowledgements
We would like to thank Hacks/Hackers for advertising the study and helping with the recruitment. We also thank members of the UW Social Futures Lab and Behavioral Data Science Group for their suggestions and feedback. This work was supported in part by the National Science Foundation’s Convergence Accelerator program under Award No. 49100421C0037.
Author contributions
X.Z., A.X.Z., and T.A. designed the study. X.Z. led, and A.S. assisted in developing the model. A.S. built the web interface for evaluation. X.Z. prepared and analyzed the data. All authors contributed to writing the manuscript. A.X.Z. and T.A. supervised the study.
Competing interests
The authors declare no competing interests.
Additional information
Supplementary Information is available for this paper.
Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
Supplementary Materials
List of supplementary materials
Table S1
Figures S1 to S27
Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
Category
Definition
# of sources
Factuality:
- Very high:
The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces.
118
- High:
The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks.
2,313
- Mostly factual:
The source is generally accurate but may have a few uncorrected fact-check failures. It can fail up to three op-ed fact checks, especially if it is a low-volume site. While it may use biased sources occasionally, it mostly links to factual content. It is usually pro-science but may sometimes use misleading wording or offer alternative viewpoints. It is reasonably transparent and trustworthy most of the time, but caution is advised.
326
- Mixed:
The source may rely on improper sourcing or link to other biased or mixed-factual sources. It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details. Sources rejecting established scientific consensus on issues such as climate change or vaccines will receive this rating or lower. Sources identified as overt propaganda or designated as hate groups by reputable third-party evaluators will receive this rating or lower due to their inherent bias and potential spread of misleading information.
1,437
- Low:
The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda.
677
- Very low:
The source is almost always unreliable and should always be fact-checked for intentional misinformation.
252
Total: 5,123
Bias:
- Least biased:
The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced.
1,054
- Left-center:
The source has a slight to moderate liberal bias. It often publishes factual information that utilizes loaded words to favor liberal causes. It is generally trustworthy for information but may require further investigation.
850
- Right-center:
Similar to the definition of left-center bias but replacing liberal with conservative.
492
- (Extremely) left:
The source is moderately to strongly biased toward liberal causes through story selection or political affiliation. It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes. It may be untrustworthy.
402
- (Extremely) right:
Similar to the definition of (extremely) left bias but replacing liberal with conservative.
314
- Pro-science:
The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing. Legitimate science follows the scientific method, is unbiased, and does not use emotional words. The source also respects the consensus of experts in the given scientific field and strives to publish peer-reviewed science. It may have a slight political bias but adheres to scientific principles.
189
- Conspiracy- pseudoscience:
The source may publish unverifiable information not always supported by evidence. It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it.
433
- Questionable:
The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence). It may be very untrustworthy and should be fact-checked on a per-article basis.
1,390
- Satire:
The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues. It does not attempt to deceive.
148
Total: 5,272
Figure S1: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each need to be corrected, whereas generating one query may overlook some misinformed, inaccurate, or misleading claims. Bold text: the misinformed, inaccurate, or misleading claims overlooked when generating one query but captured when generating more than one query.
Figure S2: Examples that show how retrieved web pages with relatively low relevance to misinformation can promote LLM (in this case, GPT-4) hallucinations when generating corrections.
Figure S3: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
Figure S4: An example of a high-helpfulness correction from Community Notes displayed on the corresponding tweet on X and visible to the public.
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S5: Expert evaluation results for text misinformation (#=155).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S6: Expert evaluation results for multimodal (textual and visual) misinformation (#=77).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S7: Expert evaluation results for political and international-affair misinformation (#=80).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S8: Expert evaluation results for economic and business misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S9: Expert evaluation results for crime and law misinformation (#=38).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S10: Expert evaluation results for social-issue and human-right misinformation (#=30).
(a) Quality of corrections in identifying and explaining inaccuracies.
(b) Quality of corrections in generated text.
(c) Quality of corrections in references.
Figure S11: Expert evaluation results for health and medicine misinformation (#=24).
Figure S12: Images as examples used for informative image captioning with in-context learning.
1Describe an image in an informative way. Your description should be only based on the given {short caption}, {name of each person}, and {raw text}. If the image is from social media, you should start with "A screenshot of". If the image is a quote from someone, you should start with "A quote from" followed by this person’s name if there is any, then by the quoted text. If the image is an article, you should start with "An article". If the image is a photo, you should start with "A photo of". If the image is a map, you should start with "A map of". {raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother! 1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google. Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession. finance.yahoo.com Could the stock market power through a recession? ’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession? ’This would be rare.’" with a picture of a moose. The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup. It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other. It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter? Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies. The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region. It is from Aviation Week.}34short caption: {[IMAGE_CAPTION]}35name of each person: {[CELEBRITIES]}36raw text: {[OCR]}37image description:
Figure S13: LLM prompt for informative image captioning.
Figure S14: Distribution of helpfulness scores of laypeople’s corrections in X Community Notes. All: All laypeople’s corrections in Community Notes. Average helpfulness: Laypeople’s corrections in Community Notes identified with average helpfulness and used in our study. High helpfulness: Laypeople’s corrections in Community Notes identified with high helpfulness and used in our study. For $x\pm y$, $x$: mean, $y$: standard deviation. Community Notes data are regularly updated; ours are up until February 12, 2023.
Figure S15: Distribution of creation times of laypeople’s corrections in X Community Notes used in our study. Median of the creation time of laypeople’s average-helpfulness corrections: 16 hours after the misinformation was posted. Median of the creation time of laypeople’s high-helpfulness corrections: 13 hours after the misinformation was posted.
Figure S16: Annotation instructions (page 1/7, continued on the next page).
Figure S17: Annotation instructions (page 2/7, continued on the next page).
Figure S18: Annotation instructions (page 3/7, continued on the next page).
Figure S19: Annotation instructions (page 4/7, continued on the next page).
Figure S20: Annotation instructions (page 5/7, continued on the next page).
Figure S21: Annotation instructions (page 6/7, continued on the next page).
Figure S22: Annotation instructions (page 7/7).
Figure S23: Annotation task page.
Figure S24: Impact of starting times of correcting misinformation on Muse’s performance. The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness correction was created (median: 16 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness correction was created (median: 13 hours after the corresponding misinformation was posted; Supplementary Fig. S15). The simulated starting time for Muse: The post time of the corresponding misinformation (i.e., 0 hours after the corresponding misinformation was posted).
Figure S25: Distribution of post times of misinformation in X Community Notes included in our analyses.
Figure S26: Impact of post times of misinformation on Muse’s performance.
(a) Overall quality of corrections.
(b) Quality of corrections in identifying and explaining inaccuracies.
(c) Quality of corrections in generated text.
(d) Quality of corrections in references.
[…]

### v2

@@
 Abstract
-Misinformation undermines public trust in science and democracy, particularly on social media where inaccuracies can spread rapidly.
-Experts and laypeople have shown to be effective in correcting misinformation by manually identifying and explaining inaccuracies.
-Nevertheless, this approach is difficult to scale, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
-LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce plausible but false content and references, and limitations in addressing multimodal information.
-To address these issues, we propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information.
-By retrieving contextual evidence and refutations, Muse can provide accurate and trustworthy explanations and references.
-It also describes visuals and conducts multimodal searches for correcting multimodal misinformation.
-We recruit fact-checking and journalism experts to evaluate corrections to real social media posts across 13 dimensions, ranging from the factuality of explanation to the relevance of references.
-The results demonstrate Muse’s ability to correct misinformation promptly after appearing on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality corrections from laypeople by 29%.
-This work underscores the potential of LLMs to combat real-world misinformation effectively and efficiently.
+Real-world misinformation can be partially correct and even factual but misleading.
+It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly.
+High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs.
+Despite the wide acceptance of manual correction, this is difficult promptly correct newly created misinformation and to scale this approach, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
+LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information.
+We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information.
+By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references.
+It also describes images and conducts multimodal searches to verify and correct multimodal content.
+Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts, that may or may not be misleading.
+We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
+The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%.
+This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
 Introduction
-Misinformation—including misinformed, inaccurate, and misleading information—has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5].
+Misinformation, which broadly includes partially incorrect and factual but misleading content[1, 2], has far-reaching and detrimental effects on individuals and society[3, 1, 2, 4, 5].
 It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 2, 7].
@@
 Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10].
-The result is tragic; an analysis showed that COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S.
+The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S.
 alone[11].
 Concerns over misinformation on social media have been particularly significant[12, 13, 14], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[15, 16].
-Fortunately, high-quality and prompt misinformation correction, which broadly includes identifying what part(s) of content is (or is not) misinformation and explaining why that part of the content is misinformed, inaccurate, or misleading, has effectively reduced people’s false beliefs[17, 4, 18].
-To this end, domain experts (e.g., on FactCheck.org) have been instrumental in providing accurate, detailed, and well-written explanations of whether and why a claim can be inaccurate with references.
-Subsequently, other research has shown that groups of laypeople can also identify and produce effective explanations to correct misinformation if their assessments are strategically aggregated[18, 19, 20].
-Nevertheless, combating misinformation in the digital and generative artificial intelligence (AI) era cannot solely rely on human intelligence[21, 22].
-The number of active fact-checking sites has remained stagnant since 2022[23].
-Even though crowdsourcing is more scalable than expert-led activities, keeping pace with millions of posts daily on social media is impossible.
-As Brandolini’s law indicates[24], correcting misinformation is laborious—often requiring one to search and evaluate the credibility of related articles with evidence completing or refuting the claims in the given piece and write responses as justifications—whereas creating and spreading misinformation is easy.
-The absence of sufficient coverage markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[25].
-While generative AI models such as LLMs (e.g., OpenAI’s GPT series[26]) raise concerns that they facilitate the creation of misinformation, they also potentially make scaling up the correction of misinformation possible.
+Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19].
+While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible.
+As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring one to search for related and trusted articles and write justifications—whereas creating and spreading misinformation is easy.
+As a consequence, a 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supplementary Fig.
+S1).
+Even high-quality responses suffer from limited effectiveness, when they are created after rather than before initial bursts of attention to misinformation content[4, 14].
+The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
+While generative AI models such as LLMs (e.g., OpenAI’s GPT series) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible.
 Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand.
@@
 A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[30, 31, 32, 33, 34, 35, 36, 37].
-However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating inaccuracies and backfiring (i.e., reinforcing rather than reducing false beliefs)[34, 41, 42].
+However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[31, 38, 39].
 These LLMs, as well as MisinfoCorrect[28], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
-Finally, understanding the quality of a correction necessitates a comprehensive human evaluation[30].
-However, prior uses of LLMs for misinformation correction were evaluated by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[31])[40], which is insufficient to comprehensively understand their performance in identifying and explaining inaccuracies, generated text, and references.
+Finally, understanding the quality of a correction necessitates a comprehensive evaluation of fact-checking experts due to the complexity of real-world misinformation[26].
+However, prior evaluation of LLMs was conducted by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[28])[37], which is insufficient to comprehensively understand their performance in identifying and explaining (in)accuracies, generated text, and references.
 In this article, we propose Muse, a scalable approach for multimodal misinformation correction.
-Muse makes use of an LLM and augments it with the ability to handle visuals, access timely, factual, and credible knowledge on the web, retrieve evidence that refutes or contextualizes misinformation, and generate clear explanations with accurate and trustworthy references.
-We recruited experts in fact-checking and journalism to comprehensively evaluate corrections made by Muse, and compare them to baselines including GPT-4 and high-quality corrections written by the collective efforts of laypeople on social media.
-Our assessment measures the overall quality of corrections, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness of corrections, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references.
-We find that overall Muse outperforms GPT-4 by 37% and high-quality corrections from laypeople by 29%, demonstrating Muse’s ability to correct both textual and visual misinformation across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine, promptly after appearing on social media.
+Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references.
+Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople.
+Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references.
+We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in effectively and promptly responding to potential misinformation.
+Results demonstrate Muse’s advance when the content is textual with or without images across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
 Approach
-Muse is designed to automatically generate a correction to a given piece of social media content that is potentially misinformation.
-As the input, the content can contain text with or without visual information.
-As the output, the correction should identify what part(s) of the content is (or is not) misinformation, explain why that part of the content is misinformation, and provide links as references.
-The general pipeline of Muse is illustrated in Fig.
+Muse is designed to automatically respond to content that potentially is misinformation.
+In other words, the content might be inaccurate, partially accurate, or even factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation.
+The content can contain text with or without visual information.
+The response should identify what part(s) of the content is (in)accurate, explain why that part of the content is (in)accurate, and provide links as references.
+The pipeline of Muse is illustrated in Fig.
 1.
-We start by introducing the details of Muse with a piece of text-only misinformation as the input.
-First, Muse generates queries based on an LLM from the misinformation (Fig.
+We start by introducing the details of Muse with a piece of text-only potential misinformation as the input.
+First, Muse generates queries based on an LLM from the potential misinformation (Fig.
 1b; Methods).
@@
 1b; Methods).
-After scraping the content from these web links, Muse calculates their direct relevance to the misinformation and removes irrelevant web pages (Fig.
+After scraping the content from these web links, Muse calculates their direct relevance to the potential misinformation and removes irrelevant web pages (Fig.
 1b; Methods).
@@
 1c; Methods).
-Next, Muse leverages an LLM to extract sentences or paragraphs from each of the web pages as evidence, i.e., the text either explicitly refutes the misinformation or provides additional context (Fig.
+Next, Muse leverages an LLM to extract text from each of the web pages as evidence.
+Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig.
 1d; Methods).
-Finally, Muse generates a correction to the misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig.
+Finally, Muse generates a response to the potential misinformation by providing an LLM with the extracted pieces of evidence and their source web links (Fig.
 1d; Methods).
-Note that misinformation, especially when it comes to social media, often contains extraneous information that is irrelevant to a correction, including unverifiable opinions or emojis, such as the textual content of the misinformation in Fig.
+Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig.
 1.
-Therefore, generating queries instead of simply using the text of the post improves web searches as a way of denoising misinformation; see example queries in Fig.
+Therefore, generating queries instead of simply using the post content improves web searches as a way of denoising the post content; see example queries in Fig.
 1b.
-Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs to be corrected, whereas generating one query may overlook some misinformed claims (Supplementary Fig.
-S1).
-Another concern may arise from filtering retrieved web pages based on how relevant the content is to the misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction.
+Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supplementary Fig.
+S2).
+Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction.
 However, the increase in selected web pages drastically elevates the expense of Muse (Methods).
-We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating corrections (Supplementary Fig.
-S2).
+We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supplementary Fig.
+S3).
 Moreover, as illustrated in Fig.
-1c, Muse ranks the selected web pages by their publishers’ factuality and bias.
-It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation).
-When the input is a piece of multimodal misinformation with textual and visual information, Muse first describes each image in text (Fig.
-1a) so that the misinformation can be handled by any LLM on downstream tasks, including query generation (Fig.
+1c, Muse filters and ranks the selected web pages by their publishers’ factuality and bias.
+It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
+When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig.
+1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig.
 1b), evidence extraction (Fig.
-1d), and correction generation (Fig.
+1d), and response generation (Fig.
 1d).
@@
 1a; Methods).
-Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate corrections[43].
+Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate verification and corrections[40].
 For example, even a state-of-the-art image captioning model[41] may describe the visual misinformation in Fig.
-1 as simply “list of banned books in Florida”.
-The description overlooks the listed titles of books that are essential for the visual misinformation’s correction (see more examples in Supplementary Fig.
-S3).
+1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supplementary Fig.
+S4).
 In addition, Muse conducts multimodal search on the web and computes multimodal relevance to filter out irrelevant web pages (Fig.
@@
 Muse’s evaluation was based on X Community Notes data.
-Community Notes (communitynotes.twitter.com, formerly Twitter Birdwatch) was launched in 2021 to empower people on X to collaboratively add context to potentially misleading posts.
-It has been shown to effectively reduce the spread of misinformation[20].
-Community Notes operates by enabling people on X, often laypeople, to identify whether a tweet is potentially misleading, write free-response corrections of any tweet, and assess whether other people’s corrections are helpful or unhelpful.
-The platform strategically aggregates people’s assessments with a transparent algorithm such that a correction has a higher helpfulness score if more people with diverse backgrounds (e.g., with different political ideologies) agree that the correction was helpful[20].
-Besides canceling out inaccurate judgments made by some individuals, the algorithm has been shown to empirically address the “tyranny of the majority” and coordinated attempts at gaming[20, 19].
-A correction with a sufficiently high helpfulness score is then displayed on the corresponding tweet on X and visible to the public[20] (Methods; Supplementary Fig.
-S4).
-We selected all potentially misleading tweets from Community Notes that have at least two corrections, where one correction has a high helpfulness score and the other has an average helpfulness score (Methods).
-We further generated corrections to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods).
-Experts in fact-checking and journalism were recruited, onboarded, and asked to evaluate the quality of corrections by various approaches to the same tweet (Methods).
-Experts were blinded to which approach had generated each correction.
-The evaluation contains 13 specific criteria, covering how well the correction identifies and explains inaccuracies, the quality of generated text, and the quality of references (Methods).
-It also contains the overall quality of correction by taking all 13 evaluation criteria into account (Methods).
-Our primary finding is that the overall quality of Muse-generated corrections is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig.
+Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has shown to reduce the spread of misinformation[17].
+Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies.
+A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig.
+S6).
+We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods).
+Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods).
+We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods).
+Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response.
+The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods).
+It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
+Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
 2a).
-The overall quality of Muse-generated corrections has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness corrections (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated corrections (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness corrections (mean: 5.2; $p=5\times 10^{-81}$; $N=462$).
-Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness corrections and GPT-4-generated corrections ($p=0.4$; $N=464$), the overall quality of GPT-4-generated corrections has the highest variability, and GPT-4 generates more corrections with extremely low quality.
-The standard deviation of the overall quality of GPT-4-generated corrections is 2.7, vs only 2.0 for Muse and laypeople.
-10% of GPT-4’s generated corrections have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness corrections, 3% for laypeople’s high-helpfulness corrections, and 2% for Muse-generated corrections.
-Note that laypeople’s corrections were created on average 14 hours after the tweet was posted on social media.
+The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$).
+Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness responses and GPT-4-generated responses ($p=0.4$; $N=464$), the overall quality of GPT-4-generated responses has the highest variability, and GPT-4 generates more responses with extremely low quality.
+The standard deviation of the overall quality of GPT-4-generated responses is 2.7, vs only 2.0 for Muse and laypeople.
+10% of GPT-4’s generated responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse-generated responses.
+Note that laypeople’s responses were created on average 14 hours after the tweet was posted on social media.
 Here, Muse only retrieved web pages published before the tweet was posted (Methods).
-Examining specific components of correction quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in identifying and explaining inaccuracies (Fig.
+Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig.
 2b-f).
-Experts assessed that Muse-generated corrections more explicitly identify and explain where and why the tweets are inaccurate than GPT-4’s and laypeople’s high-helpfulness corrections (Fig.
+Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig.
 2b).
-89% of Muse’s generated corrections explicitly identify and explain inaccuracies, 16% more than GPT-4-generated corrections, 29% more than laypeople’s high-helpfulness corrections, and 43% more than laypeople’s average-helpfulness corrections (Fig.
+89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig.
 2b).
-As for identifying where the tweets are inaccurate, we found that Muse more comprehensively identifies the tweets’ inaccuracies with fewer mistakes (i.e., claiming where the tweets should be inaccurate as accurate or vice versa) than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig.
+As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig.
 2c-d).
-91% of Muse’s generated corrections have at least one accurate identification without any mistake, 11% more than laypeople’s high-helpfulness corrections, 19% more than GPT-4-generated corrections, and 26% more than laypeople’s average-helpfulness corrections (Fig.
+91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig.
 2c).
-Muse has 61% of generated corrections accurately identifying all the tweets’ inaccuracies, vs GPT-4 has 38%, laypeople who produce high-helpfulness corrections have 26%, and laypeople who produce average-helpfulness correction have 17% only (Fig.
+Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig.
 2d).
-Furthermore, Muse explains why the tweets are inaccurate more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness corrections (Fig.
+Furthermore, Muse explains (in)accuracies more precisely and informatively than GPT-4 and laypeople who produce even high-helpfulness responses (Fig.
 2e-f).
-70% of corrections by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness corrections, 47% by GPT-4, and 37% for laypeople’s average-helpfulness corrections only (Fig.
+70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig.
 2e).
-Meanwhile, the average informativeness score of Muse-generated corrections is 8.4, 18% higher than GPT-4-generated corrections, 29% higher than laypeople’s high-helpfulness corrections, and 50% higher than laypeople’s average-helpfulness corrections (Fig.
+Meanwhile, the average informativeness score of Muse-generated responses is 8.4, 18% higher than GPT-4-generated responses, 29% higher than laypeople’s high-helpfulness responses, and 50% higher than laypeople’s average-helpfulness responses (Fig.
 2f).
-Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of generated text (Fig.
+Results also demonstrate that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of generated text (Fig.
 2g-k).
-Muse, which can augment GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals, exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text in correcting misinformation compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$).
-Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness corrections by laypeople and additionally less toxic than the text of average-helpfulness corrections by laypeople ($p=4\times 10^{-12}$; $N=462$).
-In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness corrections by laypeople, and 43% higher than the text of average-helpfulness corrections by laypeople (Fig.
+Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$).
+Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
+In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig.
 2g).
-74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness corrections by laypeople and 45% for GPT-4-generated text (Fig.
+74% of Muse-generated text is completely factual, vs 59% for the text of even high-helpfulness responses by laypeople and 45% for GPT-4-generated text (Fig.
 2h).
-Muse has 99% generated text without any mistake in the use of English, vs laypeople who produce even high-helpfulness corrections have 89% such text (Fig.
-2i).
-Around 91% of Muse-generated text is highly coherent and logical, vs 76% and 61% for the text of high- and average-helpfulness corrections by laypeople (Fig.
+Almost all of Muse-generated text does not have any mistake in the use of English (Fig.
+2i) and is not biased, impolite, and provoking (Fig.
+2k), and 91% is highly coherent and logical (vs 76% and 61% for laypeople, Fig.
 2j).
-Around 99% of Muse-generated text is not biased, impolite, and provoking, 12% higher than that of average-helpfulness corrections by laypeople (Fig.
-2k).
-Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness corrections in the quality of references (Fig.
+Additionally, results reveal that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in the quality of references (Fig.
 2l-n).
@@
 49% of its links result in “page-not-found” errors (Fig.
-2l), and only 76% of reachable links support or are relevant to the generated text (Fig.
+2l), and only 76% of reachable links are relevant to the generated text (Fig.
 2m).
@@
 2m).
-Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness corrections by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig.
+Meanwhile, Muse’s references are more credible than the references offered in even high-helpfulness responses by laypeople ($p=4\times 10^{-11}$; $N=744$; Fig.
 2n).
-Furthermore, we observed that the quality of Muse-generated corrections to text-only and multimodal (textual and visual) misinformation is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Fig.
+Furthermore, we observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content that potentially is misinformation is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
 3a; Supplementary Fig.
-S5-S6).
-The quality of Muse-generated corrections to text-only tweets is 29% higher than GPT-4-generated corrections ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness corrections ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness corrections ($p=10^{-67}$; $N=310$).
-The quality of Muse-generated corrections to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness corrections ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness corrections ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
-Finally, we found that the quality of Muse-generated corrections to misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than corrections by GPT-4 and even high-helpfulness corrections by laypeople (Methods; Fig.
+S7-S8).
+The quality of Muse-generated responses to text-based tweets is 29% higher than GPT-4-generated responses ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness responses ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness responses ($p=10^{-67}$; $N=310$).
+The quality of Muse-generated responses to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness responses ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness responses ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
+Finally, we found that the quality of Muse-generated responses to potential misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Methods; Fig.
 3b; Supplementary Fig.
-S7-S11).
-The quality of Muse-generated corrections is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated corrections.
-The quality of Muse-generated corrections is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness corrections by laypeople.
+S9-S13).
+The quality of Muse-generated responses is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated responses.
+The quality of Muse-generated responses is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness responses by laypeople.
 Discussion
-While concerns have arisen about LLMs in facilitating the creation of misinformation[21, 22], our work demonstrates the potential of LLMs to improve the online information ecosystem by correcting misinformed, inaccurate, and misleading information[2, 3].
-Correcting misinformation, particularly on social media, is complex.
-It requires comprehending multimodal content, understanding context involving emerging topics or events, accurately and comprehensively identifying stated inaccuracies that are subtly different from the facts, and writing easy-to-understand, precise, and informative explanations supported by trustworthy references.
-We propose Muse and demonstrate the high quality of its automatically generated corrections to misinformation on social media.
-Muse augments existing powerful LLMs (e.g., GPT-4[26]) with the capabilities of addressing visuals, accessing up-to-date knowledge, and finding accurate references nonparametrically (i.e., without the need to train or fine-tune a machine learning model) such that Muse can be easily and cheaply updated, especially compared to parametric models[35].
-Results further validate that addressing visual information is crucial in correcting multimodal misinformation, where GPT-4 struggles but Muse excels (Fig.
+While concerns have arisen about LLMs in facilitating the creation of misinformation[42, 43], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[1, 2].
+Real-world misinformation can combine (in)accurate, factually accurate but misleading, and unverifiable claims.
+Identifying a piece of content as misinformation without further interventions, done by many previous AI-driven approaches[44], has limited impact on reducing misinformation’s detrimental effects.
+Showing the identification results without explanations might even increase false beliefs[45].
+Correcting misinformation has been shown to reduce its spread and false beliefs, whose effectiveness is affected by quality, timeliness, and scalability[4, 22].
+However, research on scaling up and accelerating misinformation correction—which may include automation—is still at the early stage.
+Existing AI models struggle to identify and correct misinformation, especially on social media, where misinformation is not restricted to narrow domains and spreads rapidly.
+Correcting misinformation requires comprehending content that can be multimodal and the context beyond it that often involves emerging events.
+It requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content is (in)accurate with trustworthy references.
+We propose Muse and demonstrate the high quality of its automatically generated responses to social media posts that potentially are misinformation.
+Muse augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references.
+Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig.
 3a).
-Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining inaccuracies (Fig.
+Muse also exhibits significantly fewer “hallucinations” by having fewer errors in identifying and explaining (in)accuracies (Fig.
 2c,e), generating text that is more factual and relevant to the responded content (Fig.
@@
 2l-m) than GPT-4.
-We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently correcting misinformation.
+We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently responding to suspicious content.
 Muse is end-to-end and thereby simple to use.
-Meanwhile, our results reveal that Muse-generated corrections have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness corrections in correcting misinformation across modalities and domains (Fig.
+It is nonparametric, i.e., does not need training or fine-tuning an AI model, and thereby easily and cheaply updated, especially compared to parametric models[32].
+Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities and domains (Fig.
 2-3).
 Besides the highest accuracy and factuality (Fig.
-2c,e,g-h,l-m), Muse’s generated corrections show the highest readability by being the most explicit, fluent, and coherent (Fig.
+2c,e,g-h,l-m), Muse’s generated responses show the highest readability by being the most explicit, fluent, and coherent (Fig.
 2b,i-j).
-Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating corrections with the least toxic text and most credible references[31, 41, 42] (Fig.
+Muse can reduce the risk of a correction backfiring (i.e., reinforcing rather than reducing people’s false beliefs) by generating responses with the least toxic text and most credible references[28, 38, 39] (Fig.
 2k,n).
@@
 2d) and being more capable of correcting misinformation at scale.
-These high-quality corrections by Muse can generally be obtained two minutes after misinformation appears on social media (Methods).
-By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify corrections themselves (Fig 2,3f).
-Muse’s corrections cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods).
+These high-quality responses by Muse can generally be obtained two minutes after suspicious content (Methods).
+By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
+Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods).
 Our focus in designing Muse was in maximizing the quality of corrections.
-Considering the complexity of this task and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives.
-For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without explaining the inaccuracies.
+Considering the task’s complexity and our significant improvement in quality, the cost is relatively inexpensive compared to alternatives.
+For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
 This work also faces the following limitations.
-First, although Muse is capable of addressing and correcting multimodal misinformation with text and images, it cannot accept video inputs.
-Second, we focus on correcting misinformation in a single language, English, one of the most spoken languages in the world, whereas misinformation has become a global issue and exists in various high- and low-resource languages [45].
-Third, we evaluated Muse using real-world misinformation on a single social media platform, X, as its Community Notes system provides publicly accessible large-scale data and a transparent state-of-the-art aggregation algorithm, enabling us to compare Muse’s corrections with the high- and average-helpfulness corrections made by social media users.
+First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs.
+Second, we only focus on English, one of the most spoken languages in the world.
+Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has shown to reduce the spread of misinformation[17] and transparent.
 X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13].
-Fourth, experts assessed and compared Muse against one other LLM, GPT-4[26], which can also be seen as an ablation study, since Muse augments GPT-4 (Methods).
+Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods).
 GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[47, 48].
 Conclusion
-We proposed Muse, a nonparametric LLM for scaling up high-quality and prompt corrections of misinformation on social media, that can integrate textual and visual information, access timely updated knowledge, and generate easy-to-understand corrections with accurate and trustworthy references.
-Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness corrections on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of corrections overall.
-Muse also excels in correcting text and multimodal (textual and visual) misinformation and misinformation across politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
-This study demonstrates the potential of LLMs in correcting misinformation online effectively, scalably, promptly, and transparently.
+We proposed Muse, a nonparametric LLM to scale up and accelerate misinformation corrections on social media.
+Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references.
+Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness responses on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of responses overall.
+Muse also excels when the responded content is textual or multimodal (textual and visual) and when it is related to politics and international affairs, economy and business, crime and law, social issues and human rights, or health and medicine.
+This study demonstrates the potential of LLMs in responding to online misinformation effectively, scalably, promptly, and transparently.
 Acknowledgements
@@
 49100421C0037.
+T.A.
+and A.S.
+were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
 Author contributions
@@
 prepared and analyzed the data.
-All authors contributed to writing the manuscript.
+X.Z.
+drafted and all authors revised the manuscript.
 A.X.Z.
@@
 Table S1
-Figures S1 to S27
+Figures S2 to S29
 Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
@@
 Total: 5,272
-Figure S1: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each need to be corrected, whereas generating one query may overlook some misinformed, inaccurate, or misleading claims.
-Bold text: the misinformed, inaccurate, or misleading claims overlooked when generating one query but captured when generating more than one query.
-Figure S2: Examples that show how retrieved web pages with relatively low relevance to misinformation can promote LLM (in this case, GPT-4) hallucinations when generating corrections.
+Figure S1: Distribution of potential misinformation in X Community Notes (as of February 2023) that received its first response (gray) or first high-quality response (orange) within a certain amount of time.
+Figure S2: Examples that show that generating multiple queries helps decompose a post, which may have multiple claims that each needs to be verified, whereas generating one query may overlook some of them and hence lead to not comprehensive identifications of (in)accuracies.
+Bold text: the verification-needed claims that are overlooked when generating one query but captured when generating more than one query.
+Figure S3: Examples that show how retrieved web pages with relatively low relevance to potential misinformation can promote LLM (in this case, GPT-4) hallucinations when generating responses.
 Figure S4: Examples of informative image captions, which augment image captions with names of visually represented celebrities and embedded text (see Methods for implementation details).
-Figure S4: An example of a high-helpfulness correction from Community Notes displayed on the corresponding tweet on X and visible to the public.
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S5: Expert evaluation results for text misinformation (#=155).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S6: Expert evaluation results for multimodal (textual and visual) misinformation (#=77).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S7: Expert evaluation results for political and international-affair misinformation (#=80).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S8: Expert evaluation results for economic and business misinformation (#=38).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S9: Expert evaluation results for crime and law misinformation (#=38).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S10: Expert evaluation results for social-issue and human-right misinformation (#=30).
-(a) Quality of corrections in identifying and explaining inaccuracies.
-(b) Quality of corrections in generated text.
-(c) Quality of corrections in references.
-Figure S11: Expert evaluation results for health and medicine misinformation (#=24).
+Figure S5: Examples of Muse-generated responses to accurate, partially accurate, and factually correct but misleading content on social media.
+Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S7: Expert evaluation results for textual posts (#=155).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S8: Expert evaluation results for multimodal (textual and visual) posts (#=77).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S9: Expert evaluation results for political and international-affair posts (#=80).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S10: Expert evaluation results for economic and business posts (#=38).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S11: Expert evaluation results for crime and law posts (#=38).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S12: Expert evaluation results for social-issue and human-right posts (#=30).
+(a) Quality of responses in identifying and explaining (in)accuracies.
+(b) Quality of responses in generated text.
+(c) Quality of responses in references.
+Figure S13: Expert evaluation results for health and medicine posts (#=24).
 Figure S14: Images as examples used for informative image captioning with in-context learning.
@@
 Figure S15: LLM prompt for informative image captioning.
-Figure S14: Distribution of helpfulness scores of laypeople’s corrections in X Community Notes.
-All: All laypeople’s corrections in Community Notes.
-Average helpfulness: Laypeople’s corrections in Community Notes identified with average helpfulness and used in our study.
-High helpfulness: Laypeople’s corrections in Community Notes identified with high helpfulness and used in our study.
+Figure S16: Distribution of helpfulness scores of laypeople’s responses in X Community Notes.
+All: All laypeople’s responses in Community Notes.
+Average helpfulness: Laypeople’s responses in Community Notes identified with average helpfulness and used in our study.
+High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study.
 For $x\pm y$, $x$: mean, $y$: standard deviation.
 Community Notes data are regularly updated; ours are up until February 12, 2023.
-Figure S15: Distribution of creation times of laypeople’s corrections in X Community Notes used in our study.
-Median of the creation time of laypeople’s average-helpfulness corrections: 16 hours after the misinformation was posted.
-Median of the creation time of laypeople’s high-helpfulness corrections: 13 hours after the misinformation was posted.
+Figure S17: Distribution of creation times of laypeople’s responses in X Community Notes used in our study.
+Median of the creation time of laypeople’s average-helpfulness responses: 16 hours after the tweet was posted.
+Median of the creation time of laypeople’s high-helpfulness responses: 13 hours after the tweet was posted.
 Figure S18: Annotation instructions (page 1/7, continued on the next page).
@@
 Figure S25: Annotation task page.
-Figure S24: Impact of starting times of correcting misinformation on Muse’s performance.
-The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness correction was created (median: 16 hours after the corresponding misinformation was posted; Supplementary Fig.
-S15).
-The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness correction was created (median: 13 hours after the corresponding misinformation was posted; Supplementary Fig.
-S15).
-The simulated starting time for Muse: The post time of the corresponding misinformation (i.e., 0 hours after the corresponding misinformation was posted).
-Figure S25: Distribution of post times of misinformation in X Community Notes included in our analyses.
-Figure S26: Impact of post times of misinformation on Muse’s performance.
-(a) Overall quality of corrections.
-(b) Quality of corrections in identifying and explaining inaccuracies.
-(c) Quality of corrections in generated text.
-(d) Quality of corrections in references.
+Figure S26: Impact of starting times of responding to tweets on Muse’s performance.
+The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig.
+S17).
+The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig.
+S17).
+The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
+Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.
+Figure S28: Impact of post times of tweets on Muse’s performance.
+(a) Overall quality of responses.
+(b) Quality of responses in identifying and explaining (in)accuracies.
+(c) Quality of responses in generated text.
+(d) Quality of responses in references.
 […]

### v3

@@
 High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs.
-Despite the wide acceptance of manual correction, this is difficult promptly correct newly created misinformation and to scale this approach, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
+Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
 LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information.
@@
 It also describes images and conducts multimodal searches to verify and correct multimodal content.
-Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts, that may or may not be misleading.
+Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading.
 We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
@@
 Muse’s evaluation was based on X Community Notes data.
-Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has shown to reduce the spread of misinformation[17].
+Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17].
 Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies.
@@
 2e).
-Meanwhile, the average informativeness score of Muse-generated responses is 8.4, 18% higher than GPT-4-generated responses, 29% higher than laypeople’s high-helpfulness responses, and 50% higher than laypeople’s average-helpfulness responses (Fig.
+Meanwhile, the average informativeness score of Muse-generated responses is 7.9, 32% higher than laypeople’s high-helpfulness responses, 36% higher than GPT-4-generated responses, and 65% higher than laypeople’s average-helpfulness responses (Fig.
 2f).
@@
 Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$).
-Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=464$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
+Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
 In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig.
@@
 Second, we only focus on English, one of the most spoken languages in the world.
-Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has shown to reduce the spread of misinformation[17] and transparent.
+Third, we evaluated Muse using real social media content on a single platform, X, as its Community Notes system has been shown to reduce the spread of misinformation[17] and transparent.
 X is also a popular social media platform, and one where more than half of users consume news regularly[46] and where misinformation has been shown to diffuse faster than the truth[13].
@@
 List of supplementary materials
-Table S1
+Table LABEL:supp:table:mbfc
 Figures S1 to S29
 Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
-Category
-Definition
-# of sources
+Category Definition # of sources
 Factuality:
-- Very high:
-The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces.
-118
-- High:
-The source is mostly factual and uses mostly credible, low-biased, or high-factual sources. It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks.
+- Very high: The source is consistently factual, relies on credible information, promptly corrects errors, and has never failed any fact checks in news reporting or opinion pieces. 118
+- High: The source is mostly factual and uses mostly credible, low-biased, or high-factual sources.
+It corrects errors quickly and has failed only one news fact check and up to two op-ed fact checks.
 2,313
-- Mostly factual:
-The source is generally accurate but may have a few uncorrected fact-check failures.
+- Mostly factual: The source is generally accurate but may have a few uncorrected fact-check failures.
 It can fail up to three op-ed fact checks, especially if it is a low-volume site.
@@
 326
-- Mixed:
-The source may rely on improper sourcing or link to other biased or mixed-factual sources.
+- Mixed: The source may rely on improper sourcing or link to other biased or mixed-factual sources.
 It often has multiple failed fact checks and does not correct false information or lacks transparency, including the absence of a disclosed mission statement or ownership details.
@@
 1,437
-- Low:
-The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda.
-677
-- Very low:
-The source is almost always unreliable and should always be fact-checked for intentional misinformation.
-252
+- Low: The source is often unreliable and should be fact-checked for fake news, conspiracy theories, and propaganda. 677
+- Very low: The source is almost always unreliable and should always be fact-checked for intentional misinformation. 252
 Total: 5,123
 Bias:
-- Least biased:
-The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes). It is factual and usually sourced.
+- Least biased: The source has minimal bias and uses very few loaded words (i.e., wording that attempts to influence an audience by using an appeal to emotion or stereotypes).
+It is factual and usually sourced.
 1,054
-- Left-center:
-The source has a slight to moderate liberal bias.
+- Left-center: The source has a slight to moderate liberal bias.
 It often publishes factual information that utilizes loaded words to favor liberal causes.
@@
 850
-- Right-center:
-Similar to the definition of left-center bias but replacing liberal with conservative.
-492
-- (Extremely) left:
-The source is moderately to strongly biased toward liberal causes through story selection or political affiliation.
+- Right-center: Similar to the definition of left-center bias but replacing liberal with conservative. 492
+- (Extremely) left: The source is moderately to strongly biased toward liberal causes through story selection or political affiliation.
 It may utilize strong loaded words, publish misleading reports, and omit reporting of information that may damage liberal causes.
@@
 402
-- (Extremely) right:
-Similar to the definition of (extremely) left bias but replacing liberal with conservative.
-314
-- Pro-science:
-The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing.
+- (Extremely) right: Similar to the definition of (extremely) left bias but replacing liberal with conservative. 314
+- Pro-science: The source consists of legitimate science or is evidence-based through the use of credible scientific sourcing.
 Legitimate science follows the scientific method, is unbiased, and does not use emotional words.
@@
 189
-- Conspiracy- pseudoscience:
-The source may publish unverifiable information not always supported by evidence.
+- Conspiracy- pseudoscience: The source may publish unverifiable information not always supported by evidence.
 It may be untrustworthy for credible or verifiable information, so fact-checking and further investigation are recommended on a per-article basis when obtaining information from it.
 433
-- Questionable:
-The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence).
+- Questionable: The source exhibits one or more of the following: extreme bias, consistent promotion of propaganda or conspiracies, poor or no sourcing of credible information, a complete lack of transparency, or fake news (i.e., the deliberate attempt to publish hoaxes or disinformation for profit or influence).
 It may be very untrustworthy and should be fact-checked on a per-article basis.
 1,390
-- Satire:
-The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues.
+- Satire: The source exclusively uses humor, irony, exaggeration, or ridicule to expose and criticize people’s stupidity or vices, particularly in the context of contemporary politics and other topical issues.
 It does not attempt to deceive.
@@
 Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.
-Figure S28: Impact of post times of tweets on Muse’s performance.
+Figure S28: Impact of post times of tweets on the performance of Muse and baselines.
 (a) Overall quality of responses.

### v4

@@
 Abstract
-Real-world misinformation can be partially correct and even factual but misleading.
-It undermines public trust in science and democracy, particularly on social media, where it can spread rapidly.
-High-quality and timely correction of misinformation that identifies and explains its (in)accuracies has been shown to effectively reduce false beliefs.
+Real-world misinformation, often multimodal, can be partially correct and factual but misleading by cherry-picking, conflating correlation with causation, and using other tactics.
+Such misinformation is severely understudied, challenging to address, and harms many social domains, including democracy, economics, justice, human rights, and public health, particularly on social media, where it can spread rapidly.
+High-quality and timely correction of misinformation that identifies and explains its inaccuracies and accuracies has been shown to effectively reduce false beliefs.
 Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
@@
 By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references.
-It also describes images and conducts multimodal searches to verify and correct multimodal content.
-Fact-checking experts evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading.
-We propose and evaluate 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
-The results demonstrate Muse’s ability to promptly write high-quality responses to potential misinformation on social media—overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%.
-This work reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
+It conducts multimodal retrieval and interprets visual content to verify and correct multimodal content.
+Given the absence of a systematic and comprehensive evaluation approach, we propose and define 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
+Then, fact-checking experts correspondingly evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading.
+The results demonstrate Muse’s ability to write high-quality responses to potential misinformation—across modalities, tactics, domains, political leanings, and for information that has not previously been fact-checked online—within minutes of its appearance on social media.
+Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%.
+Our work provides a general methodological and evaluative framework to correct misinformation at scale.
+It reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
 Introduction
-Misinformation, which broadly includes partially incorrect and factual but misleading content[1, 2], has far-reaching and detrimental effects on individuals and society[3, 1, 2, 4, 5].
+Misinformation has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5].
 It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[6, 3, 7].
@@
 Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10].
-The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 of COVID-19-associated deaths between May 2021 and September 2022 in the U.S.
+The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 COVID-19-associated deaths between May 2021 and September 2022 in the U.S.
 alone[11].
@@
 While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[19, 20, 17], keeping pace with massive social media posts is impossible.
-As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring one to search for related and trusted articles and write justifications—whereas creating and spreading misinformation is easy.
-As a consequence, a 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supplementary Fig.
+As Brandolini’s law indicates[21], correcting misinformation is laborious—often requiring domain knowledge, information and media literacy, and explanatory skills—whereas creating and spreading misinformation is easy.
+As a consequence, 88% of suspicious content on X did not receive any response, and 93% did not receive a high-quality response within the first hour (according to X/Twitter Community Notes as of February 2023; Supp.
+Fig.
 S1).
-Even high-quality responses suffer from limited effectiveness, when they are created after rather than before initial bursts of attention to misinformation content[4, 14].
+Even high-quality responses suffer from limited effectiveness when they are created after rather than before initial bursts of attention to misinformation content[4, 14].
 The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[22].
-While generative AI models such as LLMs (e.g., OpenAI’s GPT series) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible.
+While generative AI models like LLMs raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible.
 Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand.
@@
 However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[26].
-For instance, GPT-4[27] (as of March 2023) and MisinfoCorrect[28] lack access to timely updated knowledge and are thereby ill-equipped to combat misinformation on emerging topics.
+Specifically, GPT-4[27] (as of March 2023), Llama-3 (as of July 2024)[28], MisinfoCorrect[29], MADR[30], JustiLM[31], and many other LLMs lack access to constantly growing and changing knowledge.
+These models are thereby ill-equipped to combat misinformation on emerging topics.
 They either do not provide or “hallucinate” references, which can be fabricated or irrelevant[26, 32].
-A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[30, 31, 32, 33, 34, 35, 36, 37].
+A growing body of literature has focused on retrieval-augmented LLMs, which can retrieve up-to-date information from Wikipedia or the entire Internet[33, 34, 35, 36, 37, 38, 39].
 However, their retrieval does not explicitly consider the factuality and bias of retrieved sources, posing risks of generating falsehood and backfiring (i.e., reinforcing rather than reducing false beliefs)[40, 41, 42].
-These LLMs, as well as MisinfoCorrect[28], also struggle with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
-Finally, understanding the quality of a correction necessitates a comprehensive evaluation of fact-checking experts due to the complexity of real-world misinformation[26].
-However, prior evaluation of LLMs was conducted by non-experts on limited aspects of their generated responses (e.g., the general quality of responses[28])[37], which is insufficient to comprehensively understand their performance in identifying and explaining (in)accuracies, generated text, and references.
+Every above LLM also struggles with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
+Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics and across domains[2, 3].
+Such misinformation can be more difficult to recognize, prevalent, and harmful[2, 3].
+Existing LLM and other AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., MADR[30], FOLK[38], and ProgramFC[43]), within a single domain (e.g., politics[30] or COVID-19[29, 39]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[44]).
+A general AI framework to practically identify and correct misinformation is demanded but absent.
+Evaluating the quality of a correction presented in natural language is another core challenge.
+Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics such as ROUGE[31, 44].
+However, what qualifies a correction as the ground truth is unclear[29], and existing metrics struggle to capture content factuality, especially when compared to fact-checking experts[26].
+The need for an accurate and comprehensive understanding of a model’s performance in identifying and explaining (in)accuracies, generated text, and references is recognized, but a solution is still lacking.
 In this article, we propose Muse, a scalable approach for multimodal misinformation correction.
 Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references.
+Muse is a nonparametric approach and therefore easily and cheaply updated, especially compared to parametric language models.
 Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople.
-Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies, the relevance, factuality, fluency, coherence, and toxicity of generated text, and the reachability, relevance, and credibility of references.
-We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in effectively and promptly responding to potential misinformation.
-Results demonstrate Muse’s advance when the content is textual with or without images across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
+Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies; the relevance, factuality, fluency, coherence, and toxicity of generated text; and the reachability, relevance, and credibility of references.
+We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation.
+Results demonstrate Muse’s advance no matter whether the potential misinformation content is textual with or without images; fabricates a story, misinterprets or misrepresents facts, lacks context, implies false or oversimplified causation, uses loaded language, presents false or biased data, has improper analogy or equivalence, and adopts other tactics; across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine; leans liberal or conservative; or whether the content has previously fact-checked online or not.
 Approach
 Muse is designed to automatically respond to content that potentially is misinformation.
-In other words, the content might be inaccurate, partially accurate, or even factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation.
-The content can contain text with or without visual information.
-The response should identify what part(s) of the content is (in)accurate, explain why that part of the content is (in)accurate, and provide links as references.
-The pipeline of Muse is illustrated in Fig.
-1.
+In other words, the content might be (partially) inaccurate or factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation.
+The content can contain text with or without visuals.
+The response should identify what part(s) of the content is or is not accurate, explain why that part of the content is or is not accurate, and provide links as references.
+Overall, Muse features three modules to reliably and efficiently address generalized misinformation: (1) a highly parallelizable relevance- and credibility-aware evidence retriever, (2) an LLM-based evidence-assisted response generator, and (3) an image describer that enables text-only LLMs to jointly interpret multimodal signals from the input.
+Fig.
+1 illustrates Muse’s pipeline.
 We start by introducing the details of Muse with a piece of text-only potential misinformation as the input.
@@
 1b.
-Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supplementary Fig.
+Meanwhile, generating multiple queries helps decompose a post, which may have multiple claims that each needs verification or correction, whereas generating one query may overlook some of the claims and hence lead to not comprehensive identifications and explanations of (in)accuracies (Supp.
+Fig.
 S2).
@@
 However, the increase in selected web pages drastically elevates the expense of Muse (Methods).
-We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supplementary Fig.
+We also observe that retrieved web pages with relatively low relevance can increase the prevalence of hallucinations when generating responses (Supp.
+Fig.
 S3).
@@
 1a; Methods).
-Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with additional features, including identification of celebrities and embedded text—information crucial for making accurate verification and corrections[40].
+Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with features crucial for making accurate verification and corrections[45] (Methods).
 For example, even a state-of-the-art image captioning model[46] may describe the visual misinformation in Fig.
-1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supplementary Fig.
+1 as simply “list of banned books in Florida.” The description overlooks the listed titles of books that are essential for the visual content’s verification and correction (see more examples in Supp.
+Fig.
 S4).
@@
 Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies.
-A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supplementary Fig.
+A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supp.
+Fig.
 S6).
-We included the tweets from Community Notes with at least two responses (#=247); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods).
+We included the tweets from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods).
 Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods).
@@
 Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
-In particular, Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig.
+Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig.
 2g).
@@
 2n).
-Furthermore, we observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content that potentially is misinformation is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
-3a; Supplementary Fig.
-S7-S8).
-The quality of Muse-generated responses to text-based tweets is 29% higher than GPT-4-generated responses ($p=6\times 10^{-28}$; $N=310$), 33% higher than laypeople’s high-helpfulness responses ($p=10^{-44}$; $N=310$), and 65% higher than laypeople’s average-helpfulness responses ($p=10^{-67}$; $N=310$).
-The quality of Muse-generated responses to multimodal tweets is generally 21% higher than laypeople’s high-helpfulness responses ($p=3\times 10^{-8}$; $N=154$), 39% higher than laypeople’s average-helpfulness responses ($p=10^{-16}$; $N=154$), and 56% higher than GPT-4 ($p=6\times 10^{-18}$; $N=152$).
-Finally, we found that the quality of Muse-generated responses to potential misinformation across domains including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Methods; Fig.
-3b; Supplementary Fig.
-S9-S13).
-The quality of Muse-generated responses is 35% (political and international affair; $p=10^{-16}$; $N=160$), 47% (economy and business; $p=4\times 10^{-11}$; $N=76$), 44% (crime and law; $p=3\times 10^{-8}$; $N=76$), 31% (social issues and human rights; $p=0.0002$; $N=60$), and 27% (health and medicine; $p=3\times 10^{-5}$; $N=48$) higher than GPT-4-generated responses.
-The quality of Muse-generated responses is 33% (political and international affair; $p=10^{-21}$; $N=160$), 27% (economy and business; $p=10^{-8}$; $N=76$), 28% (crime and law; $p=3\times 10^{-9}$; $N=76$), 38% (social issues and human rights; $p=7\times 10^{-10}$; $N=60$), and 25% (health and medicine; $p=4\times 10^{-5}$; $N=48$) higher than even high-helpfulness responses by laypeople.
+We further analyzed the robustness of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics.
+We first observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content is consistently higher than responses by GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig.
+3a).
+GPT-4 can be comparable to laypeople who write high-helpfulness responses when responding to text-only tweets.
+However, its performance significantly declines when responding to multimodal tweets, often resulting in a quality lower than laypeople who produce average-helpfulness responses.
+Second, we removed tweets where Muse’s response references at least one related fact-checking article, accounting for 16% of all tweets.
+Notably, in only 6% of all tweets, Muse’s response exclusively references related fact-checking article(s).
+The quality of Muse-generated responses to these tweets can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch.
+We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
+3b).
+The average quality score of Muse-generated responses for these potentially more challenging tweets is 8.0, close to its 8.1 average score for all tweets.
+Third, LLMs have exhibited political bias [47], which may diminish their ability to debunk misinformation that aligns with their political stance.
+We annotated the political leaning of tweets and found that GPT-4, which has an observed liberal bias [47], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig.
+3c).
+Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3.
+This score is consistently higher than GPT-4-generated responses and even high-helpfulness responses from laypeople by at least 26%.
+Finally, we annotated the domain of tweets and tactics used to make them or part of them false or misleading.
+Results in Fig.
+3d-e indicate that the quality of Muse-generated responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%.
+Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
+Tactics primarily include fabricating, e.g., a story with or without an authentic image, or using a digitally fabricated or altered screenshot or photo; misinterpreting or misrepresenting someone’s claim, a symbol, a policy, and others; lacking context; implying false or oversimplified causation, such as conflating correlation with causation; using loaded language; presenting false, partial, or biased data, such as cherry-picking; and having improper analogy or equivalence (see examples in Supp.
+Fig.
+S7).
 Discussion
 While concerns have arisen about LLMs in facilitating the creation of misinformation[48, 49], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[2, 3].
-Real-world misinformation can combine (in)accurate, factually accurate but misleading, and unverifiable claims.
-Identifying a piece of content as misinformation without further interventions, done by many previous AI-driven approaches[44], has limited impact on reducing misinformation’s detrimental effects.
-Showing the identification results without explanations might even increase false beliefs[45].
-Correcting misinformation has been shown to reduce its spread and false beliefs, whose effectiveness is affected by quality, timeliness, and scalability[4, 22].
-However, research on scaling up and accelerating misinformation correction—which may include automation—is still at the early stage.
-Existing AI models struggle to identify and correct misinformation, especially on social media, where misinformation is not restricted to narrow domains and spreads rapidly.
-Correcting misinformation requires comprehending content that can be multimodal and the context beyond it that often involves emerging events.
-It requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content is (in)accurate with trustworthy references.
-We propose Muse and demonstrate the high quality of its automatically generated responses to social media posts that potentially are misinformation.
-Muse augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references.
+Correcting misinformation requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content (in)accurate with trustworthy references.
+Such identifications with clear explanations have been shown to reduce misinformation’s spread and people’s false beliefs, but its effectiveness is affected by quality, timeliness, and scalability[4, 22].
+However, research on scaling up and accelerating misinformation correction is still at the early stage.
+Existing AI models are ill-equipped to identify and correct misinformation on social media.
+Such misinformation often spreads rapidly and is not restricted to narrow domains.
+It can combine (in)accurate, factually accurate but misleading, and unverifiable claims.
+It can employ various tactics, deliberately or accidentally, to create falsehood or misleadingness.
+Addressing it requires comprehending content that can be multimodal and the context beyond it that often involves emerging events.
+Existing methods to evaluate the quality of corrections presented in natural language also struggle to be accurate and comprehensive.
+Such evaluation relies on professional knowledge and skills.
+It should thoroughly assess the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
+We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, and 13 dimensions of misinformation correction quality.
+Evaluation by fact-checking experts demonstrates the high quality of Muse’s automatically generated responses to social media posts that potentially are misinformation.
 Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig.
@@
 2l-m) than GPT-4.
-We provide Muse as a solution to assist social media users and platforms in accurately, scalably, promptly, and transparently responding to suspicious content.
+Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[26, 50].
+We provide Muse as a solution to assist social media users and platforms in reliably, scalably, promptly, and transparently responding to suspicious content.
 Muse is end-to-end and thereby simple to use.
-It is nonparametric, i.e., does not need training or fine-tuning an AI model, and thereby easily and cheaply updated, especially compared to parametric models[32].
-Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities and domains (Fig.
+It is nonparametric and thereby easily and cheaply updated[51].
+Meanwhile, our results reveal that Muse-generated responses have high quality in identifying and explaining inaccuracies, generated text, and provided references, significantly surpassing GPT-4 and laypeople who produce even high-helpfulness responses in correcting misinformation across modalities, domains, political leanings, and tactics and that has not even been fact-checked online (Fig.
 2-3).
@@
 2d) and being more capable of correcting misinformation at scale.
-These high-quality responses by Muse can generally be obtained two minutes after suspicious content (Methods).
+These high-quality responses by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods).
 By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
-Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (Methods).
+Muse can also inform users of the current lack of evidence, provide an accuracy nudge, and when applicable, express uncertainty (Supp.
+Fig.
+S8).
+Muse’s responses cost about 0.5 USD per social media post at the time of our evaluation, though this cost has now been reduced to 0.2 USD, as GPT-4’s price has lowered (as of February 2024; Methods).
 Our focus in designing Muse was in maximizing the quality of corrections.
@@
 For example, a crowd of laypeople can already cost about 0.9 USD[19] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
+Notably, we also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp.
+Fig.
+S9; Methods), suggesting Muse’s generalizability.
+Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021.
+Llama-3’s training data might have included some or even all tweets used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data).
+As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these tweets).
+Hence, it is unfair to compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in expert evaluation.
+Nevertheless, our qualitative analysis illustrated in Supp.
+Fig.
+S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component.
+Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s image describer module.
 This work also faces the following limitations.
@@
 X is also a popular social media platform, and one where more than half of users consume news regularly[52] and where misinformation has been shown to diffuse faster than the truth[13].
-Fourth, experts assessed and compared Muse against one other LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (Methods).
+Fourth, experts assessed and compared Muse against one foundation LLM, GPT-4, which can also be seen as an ablation study, since Muse augments GPT-4 (see Methods and Supp.
+Fig.
+S24 for additional results of our ablation study).
 GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[53, 54].
 Conclusion
-We proposed Muse, a nonparametric LLM to scale up and accelerate misinformation corrections on social media.
+We proposed Muse, a nonparametric LLM to identify and explain (in)accuracies in social media content.
 Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references.
-Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who produce high-helpfulness responses on X Community Notes in identifying and explaining inaccuracies, generating high-quality text, providing high-quality references, and producing the highest quality of responses overall.
-Muse also excels when the responded content is textual or multimodal (textual and visual) and when it is related to politics and international affairs, economy and business, crime and law, social issues and human rights, or health and medicine.
-This study demonstrates the potential of LLMs in responding to online misinformation effectively, scalably, promptly, and transparently.
+Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who write high-helpfulness responses on X Community Notes in identifying and explaining (in)accuracies, generating high-quality text, providing high-quality references, and producing high-quality responses overall.
+Muse excels when the responded content is textual or multimodal, related to a broad range of domains, false or misleading by applying various tactics, and even when the content has not been fact-checked online.
+Muse also performs consistently well when responding to liberal and conservative content.
+This work demonstrates the potential of LLMs to respond to online misinformation in a scalable, prompt, reliable, and transparent manner—addressing key bottlenecks where existing studies struggle.
+This work also contributes to a comprehensive framework, incorporating with 13 associated standards (Methods, where we analyzed their importance), for evaluating misinformation correction models.
+Additionally, we publicly release rich annotations on social media posts covering aspects such as political leanings, domains, and tactics that make them false or misleading, along with expert-annotated responses to enable research.
 Acknowledgements
@@
 Table LABEL:supp:table:mbfc
-Figures S1 to S29
+Figures S1 to S25
 Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
@@
 Figure S6: An example of a high-helpfulness response from Community Notes displayed on the corresponding tweet and visible to the public.
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S7: Expert evaluation results for textual posts (#=155).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S8: Expert evaluation results for multimodal (textual and visual) posts (#=77).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S9: Expert evaluation results for political and international-affair posts (#=80).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S10: Expert evaluation results for economic and business posts (#=38).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S11: Expert evaluation results for crime and law posts (#=38).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S12: Expert evaluation results for social-issue and human-right posts (#=30).
-(a) Quality of responses in identifying and explaining (in)accuracies.
-(b) Quality of responses in generated text.
-(c) Quality of responses in references.
-Figure S13: Expert evaluation results for health and medicine posts (#=24).
+Figure S7: Examples of social media content that uses various tactics to make it or part of it false or misleading.
+These tactics include a: Fabricating a news event (#1), a story with an authentic image (#2, also related to conspiracy theory), and an image with AI (#3).
+b: Lacking context regarding the motivation behind behavior.
+Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers.
+c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi SS runes but a shorthand for the 46th Separate Airmobile Brigade (#2).
+d: Using loaded language.
+e: Improperly analogizing or equating masks blocking respiratory droplets and underwear blocking gas molecules (#1), and COVID-19 and cancer (#2).
+f: Presenting false, partial, or biased data.
+365348 includes multiple reports for the same child and runaways who may not be considered missing because of their guardians know their whereabouts (#1).
+The data ends in 1885 and is from a specific high elevation site in Greenland, not global temperatures (#2).
+g: Implying false or oversimplified causation.
+Many COVID-19 deaths occurred among Medicare beneficiaries because Medicare primarily serves the groups who are at greater risk of adverse outcomes from COVID-19 (#1).
+The boy was shot when—but not because—he was campaigning for Raphael Warnock (#2).
+Note that one content may apply more than one tactic intentionally or accidentally.
+Figure S8: Examples of Muse-generated responses to social media posts that lack evidence to verify.
+Figure S9: Comparison between responses generated by GPT-4, Llama-3, and their based Muse.
+We observed that GPT4-based Muse and Llama3-based Muse frequently generate similar responses to social media content (see the top example).
+GPT4-based Muse occasionally generates more comprehensive corrections than Llama3-based Muse (see the middle example).
+Note that the knowledge cutoff for Llama3 is December 2023, whereas that of GPT-4 is September 2021.
+In other words, Llama-3’s training data include more recent events and knowledge than GPT-4’s (see the bottom example, where the post was posted in February 2023, and Llama-3 more accurately identifies its falsehood than GPT-4).
 Figure S10: Images as examples used for informative image captioning with in-context learning.
@@
 It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter?
-Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies.
+Percentage of tweets from each journalist that are at-replies BDUTT 64%29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies.
 The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region.
@@
 The simulated starting time for Muse (avg): Thirty minutes before the corresponding laypeople’s average-helpfulness responses was created (median: 16 hours after the corresponding tweet was posted; Supplementary Fig.
-S17).
+S13).
 The simulated starting time for Muse (high): Thirty minutes before the corresponding laypeople’s high-helpfulness responses was created (median: 13 hours after the corresponding tweet was posted; Supplementary Fig.
-S17).
+S13).
 The simulated starting time for Muse: The post time of the corresponding tweet (i.e., 0 hours after the corresponding tweet was posted).
-Figure S27: Distribution of post times of tweets in X Community Notes included in our analyses.
 Figure S23: Impact of post times of tweets on the performance of Muse and baselines.
@@
 […]
+(a) Correlation for all responses.
+(b) Correlation for responses whose overall quality is relatively low vs high.
+Figure S25: Spearman correlation coefficient between overall response quality and each criterion that specifies overall response quality.

### v5

@@
-suppReferences
+Correcting misinformation on social media with a large language model
 Abstract
-Real-world misinformation, often multimodal, can be partially correct and factual but misleading by cherry-picking, conflating correlation with causation, and using other tactics.
-Such misinformation is severely understudied, challenging to address, and harms many social domains, including democracy, economics, justice, human rights, and public health, particularly on social media, where it can spread rapidly.
-High-quality and timely correction of misinformation that identifies and explains its inaccuracies and accuracies has been shown to effectively reduce false beliefs.
-Despite the wide acceptance of manual correction, it is difficult to be timely and scalable, a concern as technologies like large language models (LLMs) make misinformation easier to produce.
-LLMs also have versatile capabilities that could accelerate misinformation correction—however, they struggle due to a lack of recent information, a tendency to produce false content, and limitations in addressing multimodal information.
-We propose Muse, an LLM augmented with access to and credibility evaluation of up-to-date information.
-By retrieving evidence as refutation or supporting context, Muse identifies and explains (in)accuracies in a piece of content—not presupposed to be misinformation—with references.
-It conducts multimodal retrieval and interprets visual content to verify and correct multimodal content.
-Given the absence of a systematic and comprehensive evaluation approach, we propose and define 13 dimensions of misinformation correction quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
-Then, fact-checking experts correspondingly evaluate responses to social media content that are not presupposed to be (non-)misinformation but broadly include incorrect, partially correct, and correct posts that may or may not be misleading.
-The results demonstrate Muse’s ability to write high-quality responses to potential misinformation—across modalities, tactics, domains, political leanings, and for information that has not previously been fact-checked online—within minutes of its appearance on social media.
-Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29%.
-Our work provides a general methodological and evaluative framework to correct misinformation at scale.
-It reveals LLMs’ potential to help combat real-world misinformation effectively and efficiently.
+Real-world information, often multimodal, can be misinformed or potentially misleading due to factual errors, outdated claims, missing context, misinterpretation, and more.
+Such “misinformation” is understudied, challenging to address, and harms many social domains—particularly on social media, where it can spread rapidly.
+Manual correction that identifies and explains its (in)accuracies is widely accepted but difficult to scale.
+While large language models (LLMs) can generate human-like language that could accelerate misinformation correction, they struggle with outdated information, hallucinations, and limited multimodal capabilities.
+We propose Muse, an LLM augmented with vision-language modeling and web retrieval over relevant, credible sources to generate responses that determine whether and which part(s) of the given content can be misinformed or potentially misleading, and to explain why with grounded references.
+We further define a comprehensive set of rubrics to measure response quality, ranging from the accuracy of identifications and factuality of explanations to the relevance and credibility of references.
+Results show that Muse consistently produces high-quality outputs across diverse social media content (e.g., modalities, domains, political leanings), including content that has not previously been fact-checked online.
+Overall, Muse outperforms GPT-4 by 37% and even high-quality responses from social media users by 29%.
+Our work provides a general methodological and evaluative framework for correcting misinformation at scale.
+Keywords
+Misinformation, large language model, retrieval-augmented generation, multimodality, social media
+Significance Statement
+Misinformed and potentially misleading content are pervasive on social media but difficult to address at scale due to multimodality, diversity, and rapidly emerging topics.
+Such content can undermine democratic processes, economic stability, public health, and national security.
+We introduce a large language model augmented with vision-language modeling and credibility-aware web retrieval, along with a comprehensive evaluation framework.
+The model identifies and explains potentially misleading claims with grounded references, enabling transparent corrections that people can understand and audit.
+We demonstrate effectiveness and generalizability on real-world social media content through evaluations with fact-checking experts and user perception studies.
 Introduction
-Misinformation has far-reaching and detrimental effects on individuals and society[1, 2, 3, 4, 5].
+Misinformation has far-reaching and detrimental effects on individuals and society[1, 11, 3, 4, 5, 6].
 It erodes public trust in government, decreases civil engagement in elections, and has been viewed as a threat to democracy[7, 3, 8].
-Evidence has demonstrated that election misinformation helped fuel the January 6th attack on the U.S.
-Capitol, where five people died and more than 100 police officers were injured[8].
 Misinformation also drastically increases during outbreaks and disasters, as seen with the “infodemic” during the COVID-19 pandemic, which significantly increased vaccine hesitancy[9, 10].
-The result is tragic, where COVID-19 vaccines could have prevented at least 232,000 COVID-19-associated deaths between May 2021 and September 2022 in the U.S.
-alone[11].
 Concerns over misinformation on social media have been particularly significant[11, 12, 13], as the social media context interferes with truth discernment, where users post content without professional moderation and often consume news in a hasty and distracted way[14, 15].
-Fortunately, high-quality and timely correction of misinformation, which identifies what part(s) of its content is or is not accurate and explains why that part of the content is (in)accurate with references, has been shown to effectively reduce the spread of misinformation and false beliefs[17, 18, 4, 19].
+Fortunately, high-quality and timely correction of misinformation—identifying what part(s) of its content is or is not accurate and explaining why, with supporting references, has been shown to effectively reduce the spread of misinformation and false beliefs[16, 17, 4, 18].
 While domain experts (e.g., on FactCheck.org) and groups of laypeople (e.g., on X Community Notes, formerly Twitter Birdwatch) have played pivotal roles in correcting misinformation[18, 19, 16], keeping pace with massive social media posts is impossible.
@@
 The absence of sufficient coverage also markedly diminishes the impact of correction and, as the implied truth effect suggests, may even increase the perceived accuracy of misinformation that escapes correction[21].
-While generative AI models like LLMs raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible.
+While generative artificial intelligence (AI) models like large language models (LLMs) raise concerns that they facilitate creating misinformation, they also potentially make scaling up and accelerating misinformation’s correction possible.
 Recent LLMs have exhibited proficiency in generating fluent and coherent text, laying a foundation for producing explanations that the public can understand.
-Indeed, LLMs have revolutionized the field of AI and presented remarkable capabilities across domains and tasks[23, 24, 25].
 However, accurate and trustworthy misinformation correction is inseparable from accessing up-to-date and reliable information, providing accurate references to back up claims, and addressing textual and visual information, all areas where existing LLMs fall short[22].
@@
 Every above LLM also struggles with counteracting multimodal misinformation due to their nonacceptance of visual inputs.
-Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics and across domains[2, 3].
+Real-world misinformation commonly encompasses content that is partially correct and even factual but misleading through various tactics (e.g., cherry-picking, conflating correlation with causation) and across domains[11, 3].
 Such misinformation can be more difficult to recognize, prevalent, and harmful[11, 3].
-Existing LLM and other AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., MADR[30], FOLK[38], and ProgramFC[43]), within a single domain (e.g., politics[30] or COVID-19[29, 39]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[44]).
+Existing AI models to tackle misinformation focus narrowly on it that is factually incorrect (e.g., SAFE[39], MADR[26], FOLK[34], ProgramFC[40]), within a single domain (e.g., politics[26] or COVID-19[25, 35]), or employing one specific tactic (e.g., repurposing authentic images with fabricated stories[41]).
 A general AI framework to practically identify and correct misinformation is demanded but absent.
-Evaluating the quality of a correction presented in natural language is another core challenge.
-Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics such as ROUGE[31, 44].
+Evaluating the quality of a correction in natural language is another core challenge.
+Related studies heavily rely on automatic evaluation, which calculates the similarity between model-generated and “ground-truth” text using metrics like ROUGE[27, 41].
 However, what qualifies a correction as the ground truth is unclear[25], and existing metrics struggle to capture content factuality, especially when compared to fact-checking experts[22].
+Furthermore, how the general public perceives and reacts to AI-generated corrections remains unmeasured.
 The need for an accurate and comprehensive understanding of a model’s performance in identifying and explaining (in)accuracies, generated text, and references is recognized, but a solution is still lacking.
@@
 Muse makes use of an LLM and augments it with the ability to handle images, access timely and credible knowledge on the web, retrieve evidence that refutes or contextualizes the given content that may or may not be misinformation, and generate clear explanations with accurate and trustworthy references.
-Muse is a nonparametric approach and therefore easily and cheaply updated, especially compared to parametric language models.
+Muse does not need to finetune any AI models and therefore easily and cheaply updated.
 Fact-checking experts comprehensively evaluate Muse-generated responses to real social media posts that potentially are misinformation, and compare them to baselines including GPT-4 and high-quality responses based on the collective efforts of laypeople.
 Our assessment measures the overall quality of a response, specifically defined as the explicitness, accuracy, comprehensiveness, and informativeness when it identifies and explains (in)accuracies; the relevance, factuality, fluency, coherence, and toxicity of generated text; and the reachability, relevance, and credibility of references.
-We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation.
-Results demonstrate Muse’s advance no matter whether the potential misinformation content is textual with or without images; fabricates a story, misinterprets or misrepresents facts, lacks context, implies false or oversimplified causation, uses loaded language, presents false or biased data, has improper analogy or equivalence, and adopts other tactics; across a broad range of domains, including politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine; leans liberal or conservative; or whether the content has previously fact-checked online or not.
+We find that Muse outperforms GPT-4 by 37% and even high-quality responses from laypeople by 29% in accurately and promptly responding to potential misinformation, exhibiting high generalizability.
+A study with 988 participants measuring end-user perceptions further demonstrates Muse’s superior performance in increasing people’s ability to correctly identify misinformation by 9.8%.
 Approach
 Muse is designed to automatically respond to content that potentially is misinformation.
-In other words, the content might be (partially) inaccurate or factually accurate but misleading, all of which are misinformation, or fully accurate as non-misinformation.
+In other words, the content might have multiple inaccuracies, be (partially) inaccurate, or be factual but potentially misleading, all of which are misinformation, or fully accurate as non-misinformation.
 The content can contain text with or without visuals.
-The response should identify what part(s) of the content is or is not accurate, explain why that part of the content is or is not accurate, and provide links as references.
-Overall, Muse features three modules to reliably and efficiently address generalized misinformation: (1) a highly parallelizable relevance- and credibility-aware evidence retriever, (2) an LLM-based evidence-assisted response generator, and (3) an image describer that enables text-only LLMs to jointly interpret multimodal signals from the input.
+The response should identify which part(s) of the content is or is not accurate, explain why, and provide reference links.
+Overall, Muse consists of three interconnected modules to address misinformation in real-world settings: (1) an LLM for response generation, (2) a hierarchical, credibility-aware web retriever for evidence extraction, and (3) a multimodal integrator that equips modules 1&2 with multimodal capabilities.
 Fig.
@@
 We start by introducing the details of Muse with a piece of text-only potential misinformation as the input.
-First, Muse generates queries based on an LLM from the potential misinformation (Fig.
+First, Muse generates queries from the potential misinformation (Fig.
 1b; Methods).
@@
 1c; Methods).
-Next, Muse leverages an LLM to extract text from each of the web pages as evidence.
+Next, Muse extracts text from each of the web pages as evidence.
 Such evidence can refute the potential misinformation, typically happening when it is misinformation with false claims, or provide additional context, which can demonstrate that the potential misinformation is accurate or part(s) of its claims are accurate (Fig.
@@
 1d; Methods).
-Note that content, especially content posted on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the textual content of the false post in Fig.
+Note that content, especially on social media, often contains extraneous information that does not need verification and is irrelevant to correction, including unverifiable opinions or emojis, such as the false post in Fig.
 1.
@@
 S2).
-Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: theoretically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction.
+Another concern may arise from filtering retrieved web pages based on how relevant the page content is to the potential misinformation: hypothetically, including all retrieved web pages increases the amount of extracted evidence, which may not hurt and perhaps even benefit correction.
 However, the increase in selected web pages drastically elevates the expense of Muse (Methods).
@@
 It starts extracting evidence from pages with the highest factuality and least bias and then continues down the ranking, stopping when it has obtained sufficient refutations (i.e., at least two web pages were found to refute the misinformation) or gone through all the credible pages.
-When the input is a piece of multimodal (textual and visual) content, Muse first generates textual description of each image (Fig.
-1a) so that the content can be handled by any LLM on downstream tasks, including query generation (Fig.
-1b), evidence extraction (Fig.
-1d), and response generation (Fig.
-1d).
-Specifically, Muse augments image captioning models developed to describe an image in natural language with recognizing celebrities and optical characters based on an LLM (Fig.
-1a; Methods).
-Compared to existing image captioning models that capture global features of images, Muse produces more informative descriptions with features crucial for making accurate verification and corrections[45] (Methods).
+When the input is a piece of textual and visual content, Muse generates textual description of each image (Fig.
+1a) so that the content can be handled by any LLM on downstream tasks, such as response generation.
+Specifically, Muse augments image captioning models with features that existing models often fail to capture[42] but are crucial for accurate verification and correction (e.g., recognition of celebrities and optical characters; Fig.1a; Methods).
 For example, even a state-of-the-art image captioning model[43] may describe the visual misinformation in Fig.
@@
 Evaluation
-Muse’s evaluation was based on X Community Notes data.
-Community Notes empowers people on X, often laypeople, to collaboratively fact-check tweets, which has been shown to reduce the spread of misinformation[17].
+Musewas evaluated on X Community Notes data.
+Community Notes empowers social media users, often laypeople, to collaboratively fact-check posts.
+It has been shown to reduce the spread of misinformation[16] and is becoming increasingly impactful as it expands to other platforms, such as Meta’s[12].
 Every laypeople’s free-response fact-check is associated with a helpfulness score by aggregating the assessments of people with diverse backgrounds, e.g., different political ideologies.
-A response with a sufficiently high helpfulness score is then displayed on the corresponding tweet and publicly visible[17] (Supp.
+A response with a sufficiently high helpfulness score is then displayed on the corresponding post and publicly visible[16] (Supp.
 Fig.
 S6).
-We included the tweets from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods).
-Though we do not presuppose the accuracy of the tweets in Muse’s design and evaluation, we found that more than half of the tweets are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods).
-We further generated responses to these tweets based on Muse and GPT-4 (as of June 2023) (Discussion; Methods).
-Experts in fact-checking and journalism evaluated the quality of responses by various approaches to the same tweet (Methods); they were blinded to which approach had generated each response.
-The evaluation contains 13 specific criteria, covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references (Methods).
-It also contains the overall quality of a response by taking all 13 evaluation criteria into account (Methods).
+We included the posts from Community Notes with at least two responses ($n=247$, posted from February 2021 to February 2023); one has a high helpfulness score and the other has an average helpfulness score (as of February 2023; Methods).
+Though we do not presuppose the accuracy of the posts in Muse’s design and evaluation, we found that more than half of the posts are not fully (in)accurate or misleading but frequently presented in a way that combines accurate claims and inaccurate or misleading claims (Methods).
+We further generated responses to these posts based on Muse and Muse’s variants, including GPT-4 (as of June 2023) (Methods).
+Next, we defined 13 specific criteria—covering how well a response identifies and explains (in)accuracies, the quality of generated text, and the quality of references—to comprehensively assess response quality (Methods).
+Given the professional knowledge required for such evaluations, we recruited experts in fact-checking and journalism to assess and compare responses generated by various approaches to the same post, using each criterion and an overall quality score that considers all 13 criteria (Methods).
+The evaluators were blinded to which approach generated each response.
 Our primary finding is that the overall quality of Muse-generated responses is higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
@@
 The overall quality of Muse-generated responses has an average score of 8.1 out of 10, 29% higher than laypeople’s high-helpfulness responses (mean: 6.3; $p=3\times 10^{-48}$, by Mann-Whitney U test unless otherwise specified; $N=464$), 37% higher than GPT-4-generated responses (mean: 5.9; $p=4\times 10^{-42}$; $N=464$), and 56% significantly higher than laypeople’s average-helpfulness responses (mean: 5.2; $p=5\times 10^{-81}$; $N=462$).
-Despite statistical insignificance between the overall quality of laypeople’s high-helpfulness responses and GPT-4-generated responses ($p=0.4$; $N=464$), the overall quality of GPT-4-generated responses has the highest variability, and GPT-4 generates more responses with extremely low quality.
-The standard deviation of the overall quality of GPT-4-generated responses is 2.7, vs only 2.0 for Muse and laypeople.
-10% of GPT-4’s generated responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse-generated responses.
-Note that laypeople’s responses were created on average 14 hours after the tweet was posted on social media.
-Here, Muse only retrieved web pages published before the tweet was posted (Methods).
+Despite statistical insignificance between the overall quality of responses from laypeople (high-helpfulness) and GPT-4 ($p=0.4$; $N=464$), the overall quality of GPT-4’s responses has the highest variability, and GPT-4 generates more responses with extremely low quality.
+The standard deviation of the overall quality of GPT-4 responses is 2.7, vs only 2.0 for Muse and laypeople.
+10% of GPT-4’s responses have a quality score of 0 (lowest) or 1 out of 10, whereas this proportion is 5% for laypeople’s average-helpfulness responses, 3% for laypeople’s high-helpfulness responses, and 2% for Muse.
+Note that laypeople’s responses were created on average 14 hours after the post was posted on social media.
+Here, Muse only retrieved web pages published before the post was posted (Methods).
 Examining specific components of response quality, results show that Muse outperforms GPT-4 and laypeople who produce even high-helpfulness responses in identifying and explaining (in)accuracies (Fig.
 2b-f).
-Experts assessed that Muse-generated responses more explicitly identify and explain where and why a tweet is (in)accurate than GPT-4’s and laypeople’s high-helpfulness responses (Fig.
+89% of Muse’s responses explicitly identify and explain (in)accuracies, 16% more than GPT-4’s, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig.
 2b).
-89% of Muse’s generated responses explicitly identify and explain (in)accuracies, 16% more than GPT-4-generated responses, 29% more than laypeople’s high-helpfulness responses, and 43% more than laypeople’s average-helpfulness responses (Fig.
-2b).
-As for identifying where a tweet is (in)accurate, we found that Muse more comprehensively identifies a tweet’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a tweet should be inaccurate as accurate or where a tweet should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig.
+As for identifying where a post is (in)accurate, we found that Muse more comprehensively identifies a post’s (in)accuracies with fewer mistakes—here, mistakes indicate falsely claiming where a post should be inaccurate as accurate or where a post should be accurate as inaccurate—than GPT-4 and laypeople who produce even high-helpfulness responses (Fig.
 2c-d).
-91% of Muse’s generated responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4-generated responses, and 26% more than laypeople’s average-helpfulness responses (Fig.
+91% of Muse’s responses have at least one correct identification without any mistake, 11% more than laypeople’s high-helpfulness responses, 19% more than GPT-4, and 26% more than laypeople’s average-helpfulness responses (Fig.
 2c).
-Muse has 61% of generated responses accurately identifying all the (in)accuracies in a tweet, vs GPT-4 has 38%, laypeople who produce high-helpfulness responses have 26%, and laypeople who produce average-helpfulness responses have 17% only (Fig.
+Muse has 61% of responses accurately identifying all the (in)accuracies in a post, vs GPT-4 has 38%, laypeople (high helpfulness) have 26%, and laypeople (average helpfulness) have 17% only (Fig.
 2d).
@@
 2e-f).
-70% of responses by Muse have fully accurate explanations, vs 55% for laypeople’s high-helpfulness responses, 47% by GPT-4, and 37% for laypeople’s average-helpfulness responses only (Fig.
+70% of responses by Muse have fully accurate explanations, vs 55% for laypeople (high helpfulness), 47% by GPT-4, and 37% for laypeople (average helpfulness) only (Fig.
 2e).
-Meanwhile, the average informativeness score of Muse-generated responses is 7.9, 32% higher than laypeople’s high-helpfulness responses, 36% higher than GPT-4-generated responses, and 65% higher than laypeople’s average-helpfulness responses (Fig.
+Meanwhile, the average informativeness score of Muse’s responses is 7.9, 32% higher than laypeople (high helpfulness), 36% higher than GPT-4, and 65% higher than laypeople (average helpfulness) (Fig.
 2f).
@@
 Muse, when it augments GPT-4 with the capabilities of accessing timely updated knowledge and addressing visuals (Methods), exhibits enhanced relevance ($p=10^{-15}$; $N=460$) and factuality ($p=2\times 10^{-20}$; $N=459$) of text compared to GPT-4 without sacrificing fluency ($p=0.6$; $N=464$), coherence ($p=0.1$; $N=459$), and toxicity ($p=0.8$; $N=464$).
-Meanwhile, Muse-generated text is more relevant to the responded tweet ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
+Meanwhile, Muse-generated text is more relevant to the responded post ($p=2\times 10^{-30}$; $N=464$), factual ($p=4\times 10^{-6}$; $N=463$), fluent ($p=2\times 10^{-10}$; $N=464$), and coherent ($p=10^{-5}$; $N=451$) than the text of high-helpfulness responses by laypeople and additionally less toxic than the text of average-helpfulness responses by laypeople ($p=4\times 10^{-12}$; $N=462$).
 Muse-generated text has an average relevance score of 8.7, 18% higher than GPT-4-generated text, 21% higher than the text of high-helpfulness responses by laypeople, and 43% higher than the text of average-helpfulness responses by laypeople (Fig.
@@
 2n).
-We further analyzed the robustness of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics.
-We first observed that the quality of Muse-generated responses to textual and multimodal (textual and visual) content is consistently higher than responses by GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig.
+We further analyzed the generalizability of Muse in addressing diverse forms of real-world misinformation, which may vary in modality, difficulty to correct, political leaning, domain, and adopted tactics.
+We first observed that the quality of Muse-generated responses to textual and multimodal (+visual) content is consistently higher than GPT-4 and even high-helpfulness responses by laypeople, with a margin of 21% or higher (Fig.
 3a).
-GPT-4 can be comparable to laypeople who write high-helpfulness responses when responding to text-only tweets.
-However, its performance significantly declines when responding to multimodal tweets, often resulting in a quality lower than laypeople who produce average-helpfulness responses.
-Second, we removed tweets where Muse’s response references at least one related fact-checking article, accounting for 16% of all tweets.
-Notably, in only 6% of all tweets, Muse’s response exclusively references related fact-checking article(s).
-The quality of Muse-generated responses to these tweets can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch.
-We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than responses by GPT-4 and even high-helpfulness responses by laypeople (Fig.
+GPT-4 can be comparable to laypeople (high helpfulness) when responding to text-only posts.
+However, its performance significantly declines when responding to multimodal posts, often resulting in a quality lower than laypeople (average helpfulness).
+Second, we removed posts where Muse’s response references at least one related fact-checking article, accounting for 16% of all posts.
+Notably, in only 6% of all posts, Muse’s response exclusively references related fact-checking article(s).
+The quality of Muse-generated responses to these posts can be more associated with Muse’s ability to summarize retrieved professional fact-checking articles, rather than fact-checking such content from scratch.
+We found that the quality of Muse-generated responses to content that has not been fact-checked online is at least 28% higher than GPT-4 and even high-helpfulness responses by laypeople (Fig.
 3b).
-The average quality score of Muse-generated responses for these potentially more challenging tweets is 8.0, close to its 8.1 average score for all tweets.
+The average quality score of Muse’s responses for these potentially more challenging posts is 8.0, close to its 8.1 average score for all posts.
 Third, LLMs have exhibited political bias [45], which may diminish their ability to debunk misinformation that aligns with their political stance.
-We annotated the political leaning of tweets and found that GPT-4, which has an observed liberal bias [47], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig.
+We annotated the political leaning of posts and found that GPT-4, which has an observed liberal bias [45], indeed performs 5% worse in correcting misinformation that aligns with liberal ideology compared to conservative ideology (Fig.
 3c).
 Nevertheless, the quality of Muse-generated responses to both liberal and conservative content has the same average score of 8.3.
-This score is consistently higher than GPT-4-generated responses and even high-helpfulness responses from laypeople by at least 26%.
-Finally, we annotated the domain of tweets and tactics used to make them or part of them false or misleading.
+This score is consistently higher than GPT-4 and even high-helpfulness responses from laypeople by at least 26%.
+Finally, we annotated the domain of posts and tactics used to make them or part of them false or misleading.
 Results in Fig.
-3d-e indicate that the quality of Muse-generated responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%.
+3d-e indicate that the quality of Muse’ responses to potential misinformation across different domains and tactics maintains an average score around 8, consistently higher than responses by GPT-4 and even high-helpfulness responses from laypeople, with a margin of at least 19%.
 Here, domains broadly include politics and international affairs, economy and business, crime and law, social issues and human rights, and health and medicine.
@@
 S7).
+We further investigated the impact of potential training data contamination and poster popularity, demonstrating that Muse’s superior performance generalizes to correcting misinformation not included in its training data, as well as misinformation posted by users with varying levels of popularity (Methods).
+We also conducted a detailed ablation study, showing the effectiveness of the proposed retrieval mechanism and multimodal integration (Methods).
+Beyond expert evaluation, we conducted an additional study with 988 laypeople to examine how social media users perceive and react to corrections generated by Muse (Methods).
+The results demonstrate Muse’s superior performance in generating corrections useful to end users.
+Notably, it is the only approach shown to significantly increase people’s capability to correctly identify misinformation, with an improvement of 9.8%, compared to even strong baselines, including GPT-4 and high-quality responses from laypeople.
 Discussion
 While concerns have arisen about LLMs in facilitating the creation of misinformation[46, 47], our work demonstrates LLMs’ potential to improve the online information ecosystem by correcting misinformation[11, 3].
-Correcting misinformation requires identifying what part(s) of the content is (in)accurate and explaining why that part of the content (in)accurate with trustworthy references.
-Such identifications with clear explanations have been shown to reduce misinformation’s spread and people’s false beliefs, but its effectiveness is affected by quality, timeliness, and scalability[4, 22].
-However, research on scaling up and accelerating misinformation correction is still at the early stage.
-Existing AI models are ill-equipped to identify and correct misinformation on social media.
-Such misinformation often spreads rapidly and is not restricted to narrow domains.
-It can combine (in)accurate, factually accurate but misleading, and unverifiable claims.
-It can employ various tactics, deliberately or accidentally, to create falsehood or misleadingness.
-Addressing it requires comprehending content that can be multimodal and the context beyond it that often involves emerging events.
-Existing methods to evaluate the quality of corrections presented in natural language also struggle to be accurate and comprehensive.
-Such evaluation relies on professional knowledge and skills.
-It should thoroughly assess the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
-We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, and 13 dimensions of misinformation correction quality.
-Evaluation by fact-checking experts demonstrates the high quality of Muse’s automatically generated responses to social media posts that potentially are misinformation.
+Correcting misinformation requires identifying which part(s) of the content are (in)accurate and explaining why with trustworthy references.
+This approach can reduce misinformation sharing and false beliefs, but effectiveness depends on the correction’s quality, timeliness, and scalability [4, 21].
+Existing approaches, still at an early stage, struggle in particular with complex, nuanced misinformation that spans domains and goes beyond straightforward factual errors, especially when addressing it requires comprehending multimodal content and situating it in rapidly evolving real-world context.
+Existing methods for evaluating the quality of natural-language corrections also struggle to be accurate and comprehensive.
+Such evaluation relies on rigorous assessment of the identification and explanation of (in)accuracies, as well as the use of references from diverse perspectives.
+Beyond content validity, evaluation should also probe likely real-world impact by examining how end users perceive and act on the correction.
+We propose Muse, which augments existing powerful LLMs (here, it is GPT-4) with the capabilities of addressing images, accessing up-to-date knowledge, and finding accurate references, as well as 13 dimensions of correction quality.
+Evaluation by fact-checking experts demonstrates the high quality of Muse-generated responses to social media posts as potential misinformation.
 Results further validate that GPT-4 struggles to effectively respond to visual content but Muse excels (Fig.
@@
 Note that addressing LLM ’hallucination’ has been a long-standing challenge stressed in many works[22, 48].
-We provide Muse as a solution to assist social media users and platforms in reliably, scalably, promptly, and transparently responding to suspicious content.
+As more and more social media platforms (e.g., Facebook, Instagram, Threads [12]) are adopting Community Notes-like approaches to address misinformation, we provide Muse as a potential solution to assist social media users in drafting responses and to help platforms directly respond to misinformation reliably, promptly, and transparently.
 Muse is end-to-end and thereby simple to use.
@@
 2d) and being more capable of correcting misinformation at scale.
-These high-quality responses by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods).
+Additionally, we empirically demonstrate that Muse-generated responses enhance people’ capability to correctly identify misinformation by 9.8% (Methods).
+These high-quality responses and effective corrections by Muse can generally be obtained within minutes of suspicious content appearing on social media (Methods).
 By transparently providing references that refer to the retrieved web pages where evidence was collected, users can become more informed and also verify responses themselves (Fig 2,3f).
@@
 For example, a crowd of laypeople can already cost about 0.9 USD[18] to identify whether a new article’s headline and lede contains misinformation without writing down the explanation.
-Notably, we also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp.
+We also changed Muse’s foundation LLM from GPT-4 to open-sourced Llama-3 (70B) and observed two Muses are comparable without modifying any other implementation details (Supp.
 Fig.
@@
 Note that the knowledge cutoff for Llama-3 is December 2023, whereas that of GPT-4 is September 2021.
-Llama-3’s training data might have included some or even all tweets used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data).
-As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these tweets).
-Hence, it is unfair to compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in expert evaluation.
+Llama-3’s training data might have included some or even all posts used in evaluation and related events, knowledge, and corrections (e.g., their Community Notes data).
+As a comparison, GPT-4’s training data might only include up to 10% (Methods, where we show that Muse’s performance remains consistent after removing these posts).
+Hence, it is unfair to directly compare between Llama-3-based and GPT-4-based Muse, and we did not include Llama-3 and its based Muse in evaluation.
 Nevertheless, our qualitative analysis illustrated in Supp.
@@
 S9 reveals the potential to significantly reduce Muse’s cost, as GPT-4 was Muse’s most costly component.
-Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s image describer module.
-This work also faces the following limitations.
-First, although Muse is capable of responding to multimodal misinformation with text and images, it cannot accept video inputs.
+Furthermore, as one of the most capable open-sourced LLMs, Llama-3 still cannot accept visual inputs (as of August 2024), emphasizing the importance of Muse’s informative image captioning module.
+This work faces several limitations.
+First, although Muse is capable of responding to misinformation with text and images, it cannot accept video inputs.
 Second, we only focus on English, one of the most spoken languages in the world.
@@
 S24 for additional results of our ablation study).
-GPT-4 was chosen as a comparison as it is one of the best performing LLMs across a wide range of tasks available today[53, 54].
-Conclusion
-We proposed Muse, a nonparametric LLM to identify and explain (in)accuracies in social media content.
-Muse can integrate textual and visual information, access timely knowledge, and generate responses that use natural languages and have accurate and trustworthy references.
-Experts assessed that Muse significantly outperforms GPT-4 and even laypeople who write high-helpfulness responses on X Community Notes in identifying and explaining (in)accuracies, generating high-quality text, providing high-quality references, and producing high-quality responses overall.
-Muse excels when the responded content is textual or multimodal, related to a broad range of domains, false or misleading by applying various tactics, and even when the content has not been fact-checked online.
-Muse also performs consistently well when responding to liberal and conservative content.
-This work demonstrates the potential of LLMs to respond to online misinformation in a scalable, prompt, reliable, and transparent manner—addressing key bottlenecks where existing studies struggle.
-This work also contributes to a comprehensive framework, incorporating with 13 associated standards (Methods, where we analyzed their importance), for evaluating misinformation correction models.
-Additionally, we publicly release rich annotations on social media posts covering aspects such as political leanings, domains, and tactics that make them false or misleading, along with expert-annotated responses to enable research.
+GPT-4 was chosen as a comparison as it is the state-of-the-art ChatGPT (as of April 2025) and one of the best performing LLMs across a wide range of tasks available today[51, 52].
 Acknowledgements
@@
 were supported in part by the Office of Naval Research (#N00014-21-1-2154), NSF IIS-1901386, and NSF CAREER IIS-2142794.
+Supplementary material
+Supplementary material is available for this paper.
 Author contributions
 X.Z., A.X.Z., and T.A.
-designed the study.
+conceptualized the work and developed the overarching research goals and aims.
 X.Z.
-led, and A.S.
-assisted in developing the model.
+led the model development.
 A.S.
-built the web interface for evaluation.
+led the assessment of end-user perceptions of corrections, and X.Z.
+led the remainder of the evaluation.
 X.Z.
-prepared and analyzed the data.
-X.Z.
-drafted and all authors revised the manuscript.
+drafted the manuscript, including visualization, and all authors revised the paper.
 A.X.Z.
@@
 supervised the study.
-Competing interests
-The authors declare no competing interests.
-Additional information
-Supplementary information is available for this paper.
-Correspondence and requests for materials should be addressed to Tim Althoff (althoff@cs.washington.edu) and Amy X. Zhang (axz@cs.uw.edu).
+Data availability
+Data, including code, used in this study are available at https://github.com/Social-Futures-Lab/MUSE.
+We comply with X Terms of Service by only releasing the IDs of tweets.
+The experts’ names are anonymized.
 Supplementary materials
@@
 Table LABEL:supp:table:mbfc
-Figures S1 to S25
+Figures S1 to S28
 Table S1: Definitions and statistics of factuality and bias categories offered by Media Bias/Fact Check (mediabiasfactcheck.com).
@@
 Here, it is the consideration that the bill contained no funding for Florida, without which could mislead readers.
-c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi SS runes but a shorthand for the 46th Separate Airmobile Brigade (#2).
+c: Misinterpreting or misrepresenting a plan, where Biden did not promise to cancel all the debt (#1), or symbol, which is not Nazi $S$ runes but a shorthand for the 46th Separate Airmobile Brigade (#2).
 d: Using loaded language.
@@
 Figure S10: Images as examples used for informative image captioning with in-context learning.
-1Describe an image in an informative way.
+   1   {internallinenumbers*} Describe an image in an informative way.
 Your description should be only based on the given {short caption}, {name of each person}, and {raw text}.
@@
 If the image is a map, you should start with "A map of".
-{raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.2short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}3name of each person: {Fran Lebowitz}4raw text: {"In real life, I assure you, there is no such thing as algebra."}5image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}6short caption: {two men in suits}7name of each person: {Jim Caviezel, Michael Emerson}8raw text: {}9image description: {A photo of Jim Caviezel and Michael Emerson in suits}10short caption: {robbie leems on twitter}11name of each person: {}12raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google.
+{raw text} may contain nonsense data that are unnecessarily included in the image description; however, {name of each person} is not, and if the concept in {raw text} has a conflict with that in {short caption} (e.g., "Robbie Lemos" versus "robbie leems" shown later), {raw text} is often the right one.
+2   {internallinenumbers*} short caption: {a woman with glasses and a quote that says, in real life, i assure you there is no such thing as algebra}    3   name of each person: {Fran Lebowitz}    4   raw text: {"In real life, I assure you, there is no such thing as algebra."}    5   image description: {A quote from Fran Lebowitz, "In real life, I assure you, there is no such thing as algebra."}    6   short caption: {two men in suits}    7   name of each person: {Jim Caviezel, Michael Emerson}    8   raw text: {}    9   image description: {A photo of Jim Caviezel and Michael Emerson in suits}    10   short caption: {robbie leems on twitter}    11   name of each person: {}    12   {internallinenumbers*} raw text: {Robbie Lemos @RobbieLemos 1d I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google.
 Just in time for #EarthDay2023, cheers brother!
-1 2 3,790}13image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google.
-Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}14short caption: {a moose}15name of each person: {}16raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession.
+1 2 3,790}    13   {internallinenumbers*} image description: {A screenshot of a post of Robbie Lemos, "I’d like to congratulate my dear friend Deep Mind on a wonderful 1st day at work today at Google.
+Just in time for #EarthDay2023, cheers brother!" The post was posted on Twitter.}    14   short caption: {a moose}    15   name of each person: {}    16   {internallinenumbers*} raw text: {Yahoo Finance @YahooFinance Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession.
 finance.yahoo.com Could the stock market power through a recession?
-’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}17image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession?
+’This would be rare.’ 09:57 22/4/2023 3.4,011 Views 1 Retweet 1 Quote 5 Likes 1 Bookmark}    17   {internallinenumbers*} image description: {A screenshot of a post from Yahoo Finance, "Typically, the stock market bottoms four to five months before a recession ends, but RBC’s research details that it has bottomed as early as nine months before the end of a recession." The post shared an article from finance.yahoo.com claiming, "Could the stock market power through a recession?
 ’This would be rare.’" with a picture of a moose.
-The post was posted at 09:57 22/4/2023.}18short caption: {a person pouring tea into a cup}19name of each person: {}20raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}21image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup.
-It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}22short caption: {two people standing next to each other with the words love is blind}23name of each person: {Nick Lachey}24raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}25image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other.
-It was written by Hayley Miller and published by MSNBC.}26short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}27name of each person: {Rahul Kanwal}28raw text: {How engaged are the most-followed journalists on Twitter?
-Percentage of tweets from each journalist that are at-replies BDUTT 64%29image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies.
-The chart was made by mattmaldre.com.}30short caption: {a graph showing the global defense budget by region}31name of each person: {}32raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}33image description: {A graph showing the global defense budget by region.
+The post was posted at 09:57 22/4/2023.}    18   short caption: {a person pouring tea into a cup}    19   name of each person: {}    20   {internallinenumbers*} raw text: {New research reveals how coffee and tea can affect risk of early death for adults with diabetes By Sandee LaMotte, CNN Updated 7:01 PM EDT, Wed April 19, 2023 f The health benefits of tea 01:10 - Source: CNN}    21   {internallinenumbers*} image description: {An article claiming, "New research reveals how coffee and tea can affect risk of early death for adults with diabetes." It attached a picture of a person pouring tea into a cup.
+It was written by Sandee LaMotte, published by CNN, and updated at 7:01 PM EDT, Wed April 19, 2023.}    22   short caption: {two people standing next to each other with the words love is blind}    23   name of each person: {Nick Lachey}    24   {internallinenumbers*} raw text: {\"Love Is Blind\" co-host faceplants with a regressive line of questioning Hayley Miller MSNBC DAILY MSNBC}    25   {internallinenumbers*} image description: {An article claiming, "’Love Is Blind’ co-host faceplants with a regressive line of questioning." It attached a picture of Nick Lachey and another person standing next to each other.
+It was written by Hayley Miller and published by MSNBC.}    26   short caption: {a bar graph that shows how engaged are the most followed journalists on twitter}    27   name of each person: {Rahul Kanwal}    28   {internallinenumbers*} raw text: {How engaged are the most-followed journalists on Twitter?
+Percentage of tweets from each journalist that are at-replies BDUTT 64% RealMikeWilbon 38% 34% ErinAndrews 31% stephenasmith 24% andersoncooper rahulkanwal 13% 13% ninagarcia 5% sardesairajdeep 1% BillSimmons maddow 1% 0% 10% 20% 30% 40% 50% 60% chart by mattmaldre.com}    29   {internallinenumbers*} image description: {A bar graph showing how engaged the most followed journalists, including Rahul Kanwal, are on Twitter through the percentage of tweets from each journalist that are at-replies.
+The chart was made by mattmaldre.com.}    30   short caption: {a graph showing the global defense budget by region}    31   name of each person: {}    32   {internallinenumbers*} raw text: {Global Defense Budgets by Region (\$ Billions) \$1,000 800 600 400 200 0 2020 2021 2022 2023 2024 2025 Asia-Pacific Latin America North America Sub-Saharan Africa Europe Middle East & North Africa Russia & Commonwealth of Independent States Source: Aviation Week}    33   image description: {A graph showing the global defense budget by region.
 It is from Aviation Week.}    34   short caption: {[IMAGE_CAPTION]}    35   name of each person: {[CELEBRITIES]}    36   raw text: {[OCR]}    37   image description:
@@
 High helpfulness: Laypeople’s responses in Community Notes identified with high helpfulness and used in our study.
-For $x\pm y$, $x$: mean, $y$: standard deviation.
+For $\pm y, xx$: mean, $y$: standard deviation.
 Community Notes data are regularly updated; ours are up until February 12, 2023.
@@
 […]
-(a) Correlation for all responses.
-(b) Correlation for responses whose overall quality is relatively low vs high.
-Figure S25: Spearman correlation coefficient between overall response quality and each criterion that specifies overall response quality.
+Figure S25: Impact of popularity of users who posted tweets on Muse’s performance.
+Figure S26: Study interface to measure end user perceptions of corrections (page 1/2, continued on the next page).
+Figure S27: Study interface to measure end user perceptions of corrections (page 2/2).
+Figure S28: We conducted a study on Prolific with 988 participants to assess the real-world impact of Muse and how people perceive and react to AI-generated responses (two-sided Student’s t-test; experiments=16).
+(a) Only Muse-generated corrections significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$).
+(b) None of the corrections made by laypeople, GPT-4, and Muse significantly impacted participants’ intention to share the misinformation, and participants’ intentions were generally low (around 2/7).
+(c) Muse, GPT-4, and high-helpfulness responses from laypeople all received significantly higher trustworthiness ratings than average-helpfulness responses from laypeople.
+(d-f) While disclosing the source of the correction (human or AI) had no significant effect on trustworthiness or intention to share, not disclosing the source significantly enhanced the correct belief that the misinformation is misleading ($p=0.0001$).